### v1

Abstract
We introduce kernel density machines (KDM), a novel density ratio estimator in a reproducing kernel Hilbert space setting. KDM applies to general probability measures on countably generated measurable spaces without restrictive assumptions on continuity, or the existence of a Lebesgue density. For computational efficiency, we incorporate a low-rank approximation with precisely controlled error that grants scalability to large-sample settings. We provide rigorous theoretical guarantees, including asymptotic consistency, a functional central limit theorem, and finite-sample error bounds, establishing a strong foundation for practical use. Empirical results based on simulated and real data demonstrate the efficacy and precision of KDM.
Keywords: density ratio estimation, reproducing kernel Hilbert space (RKHS), low-rank approximation, finite-sample guarantees
1 Introduction
Estimating the density ratio $g_{\star}\coloneqq\frac{d\mathbb{Q}}{d\mathbb{P}}$ of the probability measure $\mathbb{Q}$ with respect to $\mathbb{P}$ is central to many tasks in machine learning, statistics, and applied sciences, such as domain adaption and transfer learning (Pan and Yang, 2010), anomaly detection (Cover and Hart, 1967), causal inference (Imbens and Rubin, 2015), conditional distribution estimation (Racine, 2008; Muandet et al., 2017), and generative modeling. Extant approaches to density ratio estimation often assume that both $\mathbb{P}$ and $\mathbb{Q}$ are continuous with respect to the Lebesgue measure, restricting their applicability to fully continuous data. In contrast, we introduce a kernel-based method that applies to general probability measures on countably generated measurable spaces, encompassing settings with discrete, continuous, mixed, or even singular distributions: kernel density machines (KDM). The parsimony of KDM with respect to the underlying assumptions significantly expands the applicability of density ratio estimation.
We model $g_{\star}$ as the sum of an exogenous prior plus a function in a reproducing kernel Hilbert space (RKHS), which, while possessing attractive theoretical properties, is known to pose computational problems with large data sets. To ensure scalability in the sample size and also higher-dimensional applications, we therefore introduce a low-rank approximation that retains critical information while reducing computational costs. The approximation error associated with the low-rank density ratio is sharply controlled by the error of a low-rank approximation of the kernel matrix pertaining to the RKHS and the data points. This combination of flexibility and computational efficiency makes our method applicable to various real-world scenarios, from structured data in financial modeling, to complex conditional distributions in scientific research.
We provide rigorous theoretical results, including asymptotic consistency, a functional central limit theorem, and finite-sample bounds for our estimator, and in particular also its low-rank approximation. These guarantees make KDM reliable across different sample sizes and data configurations. As an application, we obtain statistical tests of the prior density ratio and recover the kernel two-sample test of Gretton et al. (2012) as a special case. As a second application, we obtain a novel, consistent estimator of conditional distributions. We illustrate the applicability of KDM with both simulated and real data, including tests for independence and conditional distributions. [^4] An elementary reference implementation can be downloaded from KDM.R. The empirical results suggest KDM to be useful, in particular for larger data sets and higher dimensions than usually considered in RKHS and local nonparametric kernel density estimators (Li and Racine, 2006).
Our approach builds on recent advances in kernel-based density ratio estimation, but extends beyond existing methods by applying to general probability measures. Sugiyama et al. (2012, 2008) and Filipović et al. (2024) are closest to our approach, but are restricted to Lebesgue densities and do not provide theoretical guarantees, respectively. Another related work is Zellinger et al. (2023), who provide optimal convergence rates of density ratio estimators under rather strong regularity assumptions (source conditions), as used in inverse problems and learning theory. This is in contrast to our more agnostic framework. They also do not discuss computational scalability, as we do with low-rank approximations. Another strand of literature, including (Song et al., 2008; Klebanov et al., 2021; Park and Muandet, 2020), estimates conditional expectations directly via conditional mean embeddings (CME). This is a powerful and empirically widely used technique. But there is no obvious way how to recover density ratio estimates from CMEs.
The remainder of this paper is organized as follows. Section 2 formalizes the density ratio estimation problem for general probability measures and introduces our RKHS-based estimator. Section 3 presents theoretical guarantees, including consistency and finite-sample bounds. Section 4 discusses the low-rank approximation approach and its computational advantages. In Section 5, we apply our method to empirical tasks in domain adaptation, anomaly detection, causal inference, and high-dimensional conditional distribution estimation. Section 6 illustrates the applicablity of KDM in hypothesis testing and conditional distribution estimation for both simulated and real data. Section 7 concludes and points out future directions. The appendix contains additional results and all proofs.
2 Density Ratio Learning Problem
Let ${\mathbb{P}}$ and ${\mathbb{Q}}$ be probability measures on some countably generated [^5] Assuming a countably generated $\sigma$-algebra implies that the $L^{2}$ spaces studied below are separable. See, e.g., Brezis (2011, Theorem 4.13). measurable space ${\mathcal{Z}}$, where ${\mathbb{Q}}$ is absolutely continuous with respect to ${\mathbb{P}}$,
${\mathbb{Q}}\ll{\mathbb{P}}.$ (1)
We denote the density ratio (Radon–Nikodym derivative) by $g_{\star}\coloneqq\frac{d{\mathbb{Q}}}{d{\mathbb{P}}}$.
The goal of this paper is to learn the true density ratio $g_{\star}$ under minimal assumptions from samples of ${\mathbb{P}}$ and ${\mathbb{Q}}$. To this end, we assume that
$g_{\star}\in L^{2}_{{\mathbb{P}}},$ (2)
where we write $L^{2}_{\mathbb{M}}=L^{2}({\mathcal{Z}},{\mathbb{M}})$ for the separable $L^{2}$-space of a.s.-equivalence classes of real-valued functions on ${\mathcal{Z}}$, and where ${\mathbb{M}}$ henceforth denotes a placeholder for either ${\mathbb{P}}$ or ${\mathbb{Q}}$.
As for the hypothesis space for $g_{\star}$, we assume a separable reproducing kernel Hilbert space (RKHS) ${\mathcal{H}}$ with bounded measurable reproducing kernel $k$ on ${\mathcal{Z}}$,
$\kappa_{\infty}\coloneqq\sup_{z\in{\mathcal{Z}}}k(z,z)<\infty,$ (3)
see, e.g., Steinwart and Scovel (2012); Paulsen and Raghupathi (2016) for the definition and properties of an RKHS. [^6] By the kernel property, Assumption (3) implies that $|k(z_{1},z_{2})|\leq\sqrt{k(z_{1},z_{1})k(z_{2},z_{2})}$, hence $\sup_{z_{1},z_{2}\in{\mathcal{Z}}}|k(z_{1},z_{2})|=\sup_{z\in{\mathcal{Z}}}k(z,z)$. Hence all functions $h\in{\mathcal{H}}$ are measurable and bounded, and the canonical embeddings $J_{\mathbb{M}}\colon{\mathcal{H}}\to L^{2}_{{\mathbb{M}}}$ that map $h$ onto its respective a.s.-equivalence class $J_{\mathbb{M}}h$, are Hilbert–Schmidt operators with adjoints
$J_{\mathbb{M}}^{\ast}f=\int_{{\mathcal{Z}}}k(\cdot,z)f(z)\,{\mathbb{M}}(dz),\quad f\in L^{2}_{{\mathbb{M}}},$
see, e.g., Steinwart and Scovel (2012, Section 2). Given (2), we can represent the hypothesis density ratio as
$g=p_{\star}+J_{\mathbb{P}}h,\quad h\in{\mathcal{H}},$ (4)
for some auxiliary bounded measurable prior function $p_{\star}:{\mathcal{Z}}\to{\mathbb{R}}$ such that
$\pi_{\infty}\coloneqq\sup_{z\in{\mathcal{Z}}}|p_{\star}(z)|<\infty.$ (5)
The additive splitting (4) is motivated by the fact that $g_{\star}$ itself might not be an element of ${\mathcal{H}}$ while $g_{\star}-p_{\star}$ is. To see the relevance of this approach, consider the example where ${\mathcal{Z}}={\mathbb{R}}$, $g_{\star}=1$ and $k$ is a Gaussian kernel. Indeed, this Gaussian RKHS does not contain the constant functions (in fact, it contains no polynomials, see Minh (2010)). In this example, one could thus choose the constant prior $p_{\star}=1$. In general, we neither assume that $p_{\star}$ is positive nor that it integrates to 1 with respect to ${{\mathbb{P}}}$. Hence, $p_{\star}=0$ is a possible choice. Notably, our specification of the hypothesis density ratio (4) is novel to the literature, as it is more flexible and requires only the milder assumption (2), in contrast to the stronger assumptions typically employed in previous work. For a more detailed comparison, see Section 5.2.3 below.
It remains to introduce a suitable error function ${\mathcal{E}}(h)$, measuring the distance between the true $d{\mathbb{Q}}=g_{\star}\,d{\mathbb{P}}$ and the hypothesis measure $(p_{\star}+J_{\mathbb{P}}h)\,d{\mathbb{P}}$, that can be empirically evaluated for any $h\in{\mathcal{H}}$. To this end we consider the worst-case expectation error over all normalized test functions $f\in L^{2}_{{\mathbb{P}}}$,
${\mathcal{E}}(h)$ $\coloneqq\sup_{\|f\|_{L^{2}_{{\mathbb{P}}}}\leq 1}\bigg{|}\int_{{\mathcal{Z}}}f(z)\,{\mathbb{Q}}(dz)-\int_{{\mathcal{Z}}}f(z)(p_{\star}(z)+h(z))\,{\mathbb{P}}(dz)\bigg{|}$ (6)
$=\sup_{\|f\|_{L^{2}_{{\mathbb{P}}}}\leq 1}\big{|}\langle f,g_{\star}-p_{\star}-J_{\mathbb{P}}h\rangle_{L^{2}_{{\mathbb{P}}}}\big{|}=\|g_{\star}-p_{\star}-J_{\mathbb{P}}h\|_{L^{2}_{{\mathbb{P}}}}.$
Adding Tikhonov regularization through the penalty term $\lambda\|h\|_{{\mathcal{H}}}^{2}$ for some $\lambda>0$, we arrive at the convex problem
$\underset{h\in{\mathcal{H}}}{\text{minimize}}\,\big{\{}{\mathcal{E}}(h)^{2}+\lambda\|h\|_{\mathcal{H}}^{2}\big{\}}.$ (7)
There are at least two reasons for adding the penalty term in (7). First, empirically, it helps to avoid overfitting in finite-sample estimations, such as we pursue below. Second, it asserts that problem (7) has a unique solution, which is given by
$h_{\lambda}=(J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}+\lambda)^{-1}J_{\mathbb{P}}^{\ast}(g_{\star}-p_{\star}).$ (8)
This standard result follows from, e.g., Engl et al. (1996, Theorem 5.1). In contrast, a solution to (7) does not always exist for the limit case $\lambda=0$, as the following remark clarifies.
Remark 2.1.
The operator $J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}$ on ${\mathcal{H}}$ is trace-class. It is invertible if and only if $\ker J_{\mathbb{P}}=\{0\}$ and $\dim{\mathcal{H}}<\infty$. In general, the right hand side of (8) is thus not well defined for $\lambda=0$.
The error function ${\mathcal{E}}(h)$, and the expression in (8), contain the unobservable population density ratio $g_{\star}$ and therefore are not readily available for estimations based on samples of ${\mathbb{P}}$ and ${\mathbb{Q}}$. We solve this problem by using the identity $J_{\mathbb{P}}^{\ast}g_{\star}=J_{\mathbb{Q}}^{\ast}1$ justified by the following lemma.
Lemma 2.2.
We have $J_{\mathbb{P}}^{\ast}(fg_{\star})=J_{\mathbb{Q}}^{\ast}f$ for any bounded measurable function $f$ on ${\mathcal{Z}}$.
We infer that ${\mathcal{E}}(h)^{2}=\|g_{\star}-p_{\star}\|_{L^{2}_{{\mathbb{P}}}}^{2}-2\langle J_{\mathbb{Q}}^{\ast}1-J_{\mathbb{P}}^{\ast}p_{\star},h\rangle_{\mathcal{H}}+\langle J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}h,h\rangle_{\mathcal{H}}$, and therefore problem (7) is equivalent to the bona fide convex problem
$\underset{h\in{\mathcal{H}}}{\text{minimize}}\,\big{\{}-2\langle J_{\mathbb{Q}}^{\ast}1-J_{\mathbb{P}}^{\ast}p_{\star},h\rangle_{\mathcal{H}}+\langle(J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}+\lambda)h,h\rangle_{\mathcal{H}}\big{\}},$ (9)
and its solution (8) can be represented as
$h_{\lambda}=(J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}+\lambda)^{-1}(J_{\mathbb{Q}}^{\ast}1-J_{\mathbb{P}}^{\ast}p_{\star}).$ (10)
To assess how close $p_{\star}+J_{\mathbb{P}}h_{\lambda}$ is to the true density ratio $g_{\star}$, we consider the orthogonal decomposition of the squared population error
${\mathcal{E}}(h_{\lambda})^{2}=\|g_{\star}-p_{\star}-J_{\mathbb{P}}h_{\lambda}\|_{L^{2}_{{\mathbb{P}}}}^{2}=\underbrace{\|g_{\star}-g_{0}\|_{L^{2}_{{\mathbb{P}}}}^{2}}_{\text{projection error}}+\underbrace{\|g_{0}-p_{\star}-J_{\mathbb{P}}h_{\lambda}\|_{L^{2}_{{\mathbb{P}}}}^{2}}_{\text{regularization error}},$ (11)
into the sum of the squared projection error and squared regularization error, where $g_{0}-p_{\star}$ denotes the orthogonal projection of $g_{\star}-p_{\star}$ onto the closure $\overline{\operatorname{Im}J_{\mathbb{P}}}$ of the image of $J_{\mathbb{P}}$ in $L^{2}_{{\mathbb{P}}}$. The following lemma collects some elementary facts about these error terms.
Lemma 2.3.
  1. (i)
    The projection error vanishes, $\|g_{\star}-g_{0}\|_{L^{2}_{{\mathbb{P}}}}=0$, if and only if the ground truth $g_{\star}-p_{\star}$ lies in $\overline{\operatorname{Im}J_{\mathbb{P}}}$. This holds in particular if the kernel $k$ is universal in the sense that $\overline{\operatorname{Im}J_{\mathbb{P}}}=L^{2}_{{\mathbb{P}}}$.
  2. (ii)
    The regularization error vanishes as $\lambda$ tends to zero, $\lim_{\lambda\to 0}\|g_{0}-p_{\star}-J_{\mathbb{P}}h_{\lambda}\|_{L^{2}_{{\mathbb{P}}}}=0$, albeit the convergence may be slow in general.
  3. (iii)
    The projection $g_{0}-p_{\star}=J_{\mathbb{P}}h_{0}$ is attained for some $h_{0}\in{\mathcal{H}}$ if and only if (7) admits a solution for $\lambda=0$. In this case, we have
    $\|h_{\lambda}\|_{\mathcal{H}}$ $\leq\|h_{0}\|_{\mathcal{H}},$ (12)
    $\lim_{\lambda\to 0}\|h_{\lambda}-h_{0}\|_{\mathcal{H}}$ $=0,$ (13)
    and the following rate of convergence in $\lambda$ for the regularization error holds,
    $\|g_{0}-p_{\star}-J_{\mathbb{P}}h_{\lambda}\|_{L^{2}_{{\mathbb{P}}}}\leq 2^{-1/2}\|h_{0}\|_{\mathcal{H}}\lambda^{1/2}.$ (14)
3 Sample Estimator
[…]
$S_{\mathbb{M}}h\coloneqq[h(z_{{\mathbb{M}},1}),\dots,h(z_{{\mathbb{M}},n})]^{\top},$
which are the sample analogues of $J_{\mathbb{M}}$, with adjoints $S_{\mathbb{M}}^{\ast}v=\sum_{i=1}^{n}k(\cdot,z_{{\mathbb{M}},i})v_{i}$.
Note that the sample analogue of $J_{\mathbb{M}}^{\ast}$ is $n^{-1}S_{\mathbb{M}}^{\ast}$. Hence the sample analogue of (9) is the convex problem
$\underset{h\in{\mathcal{H}}}{\text{minimize}}\,\big{\{}-2\langle S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{p}_{\star},h\rangle_{\mathcal{H}}+\langle(S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}+n\lambda)h,h\rangle_{\mathcal{H}}\big{\}},$ (15)
[…]
$\bm{p}_{\star}\coloneqq[p_{\star}(z_{{\mathbb{P}},1}),\dots,p_{\star}(z_{{\mathbb{P}},n})]^{\top}.$ (16)
The unique solution of (15) can be calculated to be
$\hat{h}_{\lambda}=(S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}+n\lambda)^{-1}\big{(}S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{p}_{\star}\big{)},$ (17)
which is our sample estimator of (10). We first provide limit theorems and finite-sample guarantees for this estimator. In the next section, we then show how to efficiently compute it.
Here is our first main result. For any bounded measurable function $g:{\mathcal{Z}}\to{\mathbb{R}}$, we define the bounded operator $\operatorname{diag}(g):L^{2}_{\mathbb{M}}\to L^{2}_{\mathbb{M}}$ by point-wise multiplication, $\operatorname{diag}(g)f\coloneqq g\cdot f$, which is a natural generalization of the vector-to-matrix $\operatorname{diag}$ operator.
Theorem 3.1.
The estimator (17) satisfies the following properties,
  1. (i)
    Asymptotic consistency: $\hat{h}_{\lambda}\to h_{\lambda}$ in ${\mathcal{H}}$ a.s. as $n\to\infty$.
  2. (ii)
    Functional central limit theorem: $n^{1/2}(\hat{h}_{\lambda}-h_{\lambda})\to{\mathcal{N}}(0,O_{\lambda})$ in distribution as $n\to\infty$, where the covariance operator $O_{\lambda}:{\mathcal{H}}\to{\mathcal{H}}$ is given by $O_{\lambda}=(J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}+\lambda)^{-1}Q_{\lambda}(J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}+\lambda)^{-1}$ for the non-negative self-adjoint trace-class operator $Q_{\lambda}:{\mathcal{H}}\to{\mathcal{H}}$ defined by
    $Q_{\lambda}$ $\coloneqq J_{\mathbb{Q}}^{\ast}J_{\mathbb{Q}}-(J_{\mathbb{Q}}^{\ast}1)\otimes(J_{\mathbb{Q}}^{\ast}1)$ (18)
    $\quad+J_{\mathbb{P}}^{\ast}\operatorname{diag}(p_{\star}+J_{\mathbb{P}}h_{\lambda})^{2}J_{\mathbb{P}}-(J_{\mathbb{P}}^{\ast}(p_{\star}+J_{\mathbb{P}}h_{\lambda}))\otimes(J_{\mathbb{P}}^{\ast}(p_{\star}+J_{\mathbb{P}}h_{\lambda})).$
  3. (iii)
    Finite-sample guarantees: for any $\eta\in(0,1)$, with sampling probability of at least $1-\eta$, we have
    $\|\hat{h}_{\lambda}-h_{\lambda}\|_{\mathcal{H}}\leq C_{FS}(\eta,\|h_{\lambda}\|_{\mathcal{H}})\,\lambda^{-1}n^{-1/2},$ (19)
    for the coefficient
    $C_{FS}(\eta,s)\coloneqq 2\sqrt{2\log(2/\eta)\kappa_{\infty}}(1+\pi_{\infty}+s\sqrt{\kappa_{\infty}}).$ (20)
4 Low-rank Approximation
In this section, we discuss efficient ways to compute the sample estimator (17). The well-known representer theorem, which follows directly from (17), states that $\hat{h}_{\lambda}$ lies in the finite-dimensional subspace $\operatorname{Im}S^{\ast}$ of ${\mathcal{H}}$, where we define the joint sampling operator by
$S:{\mathcal{H}}\to{\mathbb{R}}^{2n}\cong{\mathbb{R}}^{n}\oplus{\mathbb{R}}^{n},\quad S\coloneqq\begin{bmatrix}S_{\mathbb{P}}\\ S_{\mathbb{Q}}\end{bmatrix},$
whose adjoint is given by $S^{\ast}=\begin{bmatrix}S_{\mathbb{P}}^{\ast}&S_{\mathbb{Q}}^{\ast}\end{bmatrix}:{\mathbb{R}}^{2n}\to{\mathcal{H}}$. That is, the sample estimator can be represented by a linear combination of the kernel functions $\hat{h}_{\lambda}=\sum_{i=1}^{2n}\beta_{i}k(\cdot,z_{i})$, for some vector of coefficients $\beta\in{\mathbb{R}}^{2n}$, where we define $z_{i}\coloneqq z_{{\mathbb{P}},i}$ and $z_{n+j}\coloneqq z_{{\mathbb{Q}},j}$ for $i,j=1,\dots n$. Plugging this in (15) leads to a convex problem in $\beta$ of dimension $2n$, which in principle can be solved numerically for $\beta$. However, for large $n$, say $n\geq 10^{5}$, the computational cost can become prohibitively large.
A standard practice to tackle this problem is the Nyström method, which consists of selecting a subsample $z_{\pi_{1}},\dots,z_{\pi_{m}}$, for some pivot set $\Pi=\{\pi_{1},\dots,\pi_{m}\}\subseteq\{1,\dots,2n\}$ of size $m\leq 2n$, and approximating the kernel $k$ by projecting it on the subspace spanned by $k(\cdot,z_{\pi_{1}}),\dots,k(\cdot,z_{\pi_{m}})$ in ${\mathcal{H}}$. The Nyström approximation is well known in the machine learning literature, see, e.g., Martinsson and Tropp (2020, Section 19.2) or Chen et al. (2023). For the sake of a self-contained exposition, we nevertheless briefly review the method, thereby also providing an explicit error bound for the approximation (Lemma 4.4) that remarkably does not seem to be known in the literature. To this end, we define the symmetric positive semi-definite $(2n)\times(2n)$-kernel matrix
$\bm{K}=\begin{bmatrix}\bm{K}_{\mathbb{P}}&\bm{K}_{{\mathbb{P}}{\mathbb{Q}}}\\ \bm{K}_{{\mathbb{Q}}{\mathbb{P}}}&\bm{K}_{\mathbb{Q}}\end{bmatrix},\quad\bm{K}_{ij}\coloneqq k(z_{i},z_{j}),$ (21)
which, incidentally, is the matrix representation of the operator $SS^{\ast}$ on ${\mathbb{R}}^{2n}$. We can efficiently compute a pivot set $\Pi$ of indices via the pivoted incomplete Cholesky decomposition in Algorithm 1.888In Algorithm 1, $\bm{B}^{(0)}=[\,]$ and $\bm{L}^{(0)}=[\,]$ denote empty arrays in the sense that, in the first iteration, we have $\ell_{1}=(d^{(0)}_{\pi_{1}})^{-1/2}\bm{K}_{:,\pi_{1}}$, $b_{1}=(d^{(0)}_{\pi_{1}})^{-1/2}e_{\pi_{1}}$ and $\bm{L}^{(1)}\coloneqq\ell_{1}$, $\bm{B}^{(1)}\coloneqq b_{1}$. We denote by $e_{i}$ the standard Euclidean basis vector whose $i$-th component equals 1 and all other components are zero. For a matrix $\bm{A}$ and ordered index set $I$, we write $\bm{A}_{I,:}$, $\bm{A}_{:,I}$, for the corresponding sub-matrices.
Algorithm 1 Pivoted incomplete Cholesky decomposition
Input: kernel matrix $\bm{K}$, tolerance $\epsilon\geq 0$.
Initialize $d^{(0)}\coloneqq\operatorname{diag}\bm{K}$, $\bm{L}^{(0)}\coloneqq[\,]$, $\bm{B}^{(0)}\coloneqq[\,]$, $\Pi^{(0)}\coloneqq\emptyset$, $i\coloneqq 0$.
While $\|d^{(i)}\|_{1}>\epsilon$ (and hence $d^{(i)}\neq\bm{0}$) do the following steps:
  1. (i)
    select a pivot index $\pi_{i+1}\in\{1,\dots,2n\}\setminus\Pi^{(i)}$ such that $d^{(i)}_{\pi_{i+1}}\neq 0$;
  2. (ii)
    set $\Pi^{(i+1)}\coloneqq\Pi^{(i)}\cup\{\pi_{i+1}\}$;
  3. (iii)
    set $\ell_{i+1}\coloneqq(d^{(i)}_{\pi_{i+1}})^{-1/2}(\bm{K}_{\cdot,\pi_{i+1}}-\bm{L}^{(i)}(\bm{L}^{(i)}_{\pi_{i+1},\cdot})^{\top})$;
  4. (iv)
    set $b_{i+1}\coloneqq(d^{(i)}_{\pi_{i+1}})^{-1/2}(e_{\pi_{i+1}}-\bm{B}^{(i)}(\bm{L}^{(i)}_{\pi_{i+1},\cdot})^{\top})$;
  5. (v)
    set $\bm{L}^{(i+1)}\coloneqq[\bm{L}^{(i)},\ell_{i+1}]$, $\bm{B}^{(i+1)}\coloneqq[\bm{B}^{(i)},b_{i+1}]$;
  6. (vi)
    set $d^{(i+1)}\coloneqq d^{(i)}-\ell_{i+1}\circ\ell_{i+1}$;
  7. (vii)
    set counter $i\coloneqq i+1$;
Output: rank $m\coloneqq i$, pivot set $\Pi\coloneqq\Pi^{(m)}$, matrices $\bm{L}\coloneqq\bm{L}^{(m)}$, $\bm{R}\coloneqq\bm{B}^{(m)}_{\Pi,:}$
The following classic result asserts that Algorithm 1 works, see, e.g., Horn and Zhang (2005) or Harbrecht et al. (2012). A novelty of Algorithm 1 lies in the computation of the biorthogonal matrix $\bm{R}$ at no additional complexity along with $\bm{L}$. For completeness, we provide a self-contained proof in Appendix A.
Proposition 4.1.
[…]
$\bm{K}_{:,\Pi}\bm{R}$ $=\bm{L},$ (22)
$\bm{R}^{\top}\bm{L}_{\Pi,:}$ $=\bm{I}_{m},$ (23)
and such that $\bm{K}-\bm{L}\bm{L}^{\top}$ is positive semi-definite with trace bounded by
$\operatorname{trace}(\bm{K}-\bm{L}\bm{L}^{\top})\leq\epsilon.$ (24)
Moreover, the column vectors $\bm{K}_{\cdot,\Pi}$ are linearly independent, and the following identities hold,
$\bm{R}\bm{R}^{\top}$ $=(\bm{K}_{\Pi,\Pi})^{-1},$ (25)
$\bm{L}\bm{L}^{\top}$ $=\bm{K}_{:,\Pi}(\bm{K}_{\Pi,\Pi})^{-1}\bm{K}_{\Pi,:}.$ (26)
The alternative representation (26) reveals that $\bm{L}\bm{L}^{\top}$ is in fact the column Nyström approximation of the matrix $\bm{K}$, see, e.g., Martinsson and Tropp (2020, Section 19.2). Given the rank $m$, the computational complexity of Algorithm 1 is ${\mathcal{O}}(m^{2}n)$. However, for a given tolerance $\epsilon\geq 0$, the required rank $m$ may be large. Theoretical results with a priori bounds on $m$ are given in Chen et al. (2023), which depend on the pivoting strategy in step (i). Some of them are discussed in Appendix A.
We now turn to the approximation of $\hat{h}_{\lambda}$. The following notation is useful for this. We define the array of sample points $\bm{z}\coloneqq[z_{1},\dots,z_{2n}]^{\top}$, and similarly, $\bm{z}_{{\mathbb{P}}}$, $\bm{z}_{{\mathbb{Q}}}$, and $\bm{z}_{I}\coloneqq[z_{i_{1}},\dots,z_{i_{p}}]^{\top}$, for any ordered index set $I=\{i_{1},\dots,i_{p}\}\subseteq\{1,\dots,2n\}$. Accordingly, we denote by
$k(\cdot,\bm{z}_{I})\coloneqq[k(\cdot,z_{i_{1}}),\dots,k(\cdot,z_{i_{p}})]^{\top}$
the corresponding array of kernel functions. Similarly, for any function $h$ on ${\mathcal{Z}}$, we denote by $h(\bm{z}_{I})$ the corresponding array of function values. The following corollary is a direct consequence of Proposition 4.1, properties (22) and (23).
Corollary 4.2.
The functions $\bm{\psi}=[\psi_{1}(\cdot),\dots,\psi_{m}(\cdot)]^{\top}$ defined by
$\bm{\psi}\coloneqq\bm{R}^{\top}k(\cdot,\bm{z}_{\Pi}),$ (27)
form an orthonormal basis of the subspace ${\mathcal{H}}_{\Pi}\coloneqq\operatorname{span}\{k(\cdot,z_{\pi_{1}}),\dots,k(\cdot,z_{\pi_{m}})\}$ in ${\mathcal{H}}$. The orthogonal projection $P_{\Pi}$ in ${\mathcal{H}}$ on ${\mathcal{H}}_{\Pi}$ is for $h\in{\mathcal{H}}$ given by,
$P_{\Pi}h=\bm{\psi}^{\top}\langle\bm{\psi},h\rangle_{\mathcal{H}}=k(\cdot,\bm{z}_{\Pi}^{\top})\bm{R}\bm{R}^{\top}h(\bm{z}_{\Pi}).$ (28)
Hence $P_{\Pi}h$ can be efficiently evaluated by querying only $m$ kernel functions at a time.
We approximate the convex problem (15) over ${\mathcal{H}}$ by restricting it to the subspace ${\mathcal{H}}_{\Pi}$. This leads to the $m$-dimensional convex problem
$\underset{h\in{\mathcal{H}}_{\Pi}}{\text{minimize}}\,\big{\{}-2\langle P_{\Pi}(S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{p}_{\star}),h\rangle_{\mathcal{H}}+\langle(P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}+n\lambda)h,h\rangle_{\mathcal{H}}\big{\}}.$ (29)
The unique solution of (29) is
$\widetilde{h}_{\lambda}=(P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}+n\lambda)^{-1}P_{\Pi}(S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{p}_{\star}),$ (30)
which is the low-rank approximation of the sample estimator (17). It can be efficiently computed as the following result shows.
Lemma 4.3.
The approximated estimator (30) can be expressed in coordinates with respect to the orthonormal basis $\bm{\psi}$ as
$\widetilde{h}_{\lambda}=k(\cdot,\bm{z}_{\Pi}^{\top})\bm{R}\big{(}\bm{L}_{\mathbb{P}}^{\top}\bm{L}_{\mathbb{P}}+n\lambda\big{)}^{-1}\big{(}\bm{L}_{\mathbb{Q}}^{\top}\bm{1}-\bm{L}_{\mathbb{P}}^{\top}\bm{p}_{\star}\big{)},$ (31)
which can be computed with complexity ${\mathcal{O}}(m^{2}n)$.
The following key result provides an explicit approximation error bound.
Lemma 4.4.
The low-rank approximation error is bounded by
$\|\hat{h}_{\lambda}-\widetilde{h}_{\lambda}\|_{\mathcal{H}}\leq n^{-1/2}\lambda^{-1}C_{AE}(\epsilon,\lambda)$ (32)
for the coefficient
$C_{AE}(\epsilon,\lambda)\coloneqq\epsilon^{1/2}\big{(}1+\lambda^{-1/2}\kappa_{\infty}^{1/2}\big{)}(\pi_{\infty}+1).$ (33)
Combining the finite-sample guarantees in Theorem 3.1 (iii) with the approximation error bound in Lemma 4.4 and Lemma 2.3, we obtain finite-sample guarantees for the low-rank estimator. This is our second main result.
Theorem 4.5.
For any $\eta\in(0,1)$, with sampling probability of at least $1-\eta$, the low-rank sample estimation error is bounded by
$\|h_{\lambda}-\widetilde{h}_{\lambda}\|_{\mathcal{H}}\leq\big{(}C_{FS}(\eta,\|h_{\lambda}\|_{\mathcal{H}})+C_{AE}(\epsilon,\lambda)\big{)}\lambda^{-1}n^{-1/2},$ (34)
and the worst-case expectation error (6) by
${\mathcal{E}}(\widetilde{h}_{\lambda})\leq{\mathcal{E}}(h_{\lambda})+\kappa_{\infty}^{1/2}\big{(}C_{FS}(\eta,\|h_{\lambda}\|_{\mathcal{H}})+C_{AE}(\epsilon,\lambda)\big{)}\lambda^{-1}n^{-1/2},$ (35)
for coefficients $C_{FS}(\eta,s)$ and $C_{AE}(\epsilon,\lambda)$ given in (20) and (33). Moreover, if the ground truth
$g_{\star}-p_{\star}=J_{\mathbb{P}}h_{0}$ is attained for some $h_{0}\in{\mathcal{H}}$, (36)
then the error bound (35) can be improved as
${\mathcal{E}}(\widetilde{h}_{\lambda})\leq 2^{-1/2}\|h_{0}\|_{\mathcal{H}}\lambda^{1/2}+\kappa_{\infty}^{1/2}\big{(}C_{FS}(\eta,\|h_{0}\|_{\mathcal{H}})+C_{AE}(\epsilon,\lambda)\big{)}\lambda^{-1}n^{-1/2}.$ (37)
In view of Theorem 4.5, it is straightforward to find sequences $\lambda_{n}\to 0$ and $\epsilon_{n}\to 0$ such that $\lambda_{n}^{-1}n^{-1/2}\to 0$ and $C_{AE}(\epsilon_{n},\lambda_{n})\to 0$, so that the total low-rank estimation error bound in (37) goes to zero, as $n\to\infty$. In fact, for $\lambda_{n}=n^{-1/4}$ and $\epsilon_{n}=O(\lambda_{n})$, the right hand side of (37) is $O(n^{-1/4})$. Comparing to Li and Racine (2006, Section 2.7) for local kernel density estimators on Euclidean state spaces ${\mathcal{Z}}\subseteq{\mathbb{R}}^{d}$ of dimension $d$, their convergence rate is quoted as $O(n^{-2/(d+4)})$, with correctly specified model and optimal bandwidth. As noted in Duong and Hazelton (2005), local kernel density estimation is therefore deemed challenging for dimensions $d$ beyond four, even with large samples.
5 Applications
In this section, we illustrate two applications of KDM. The first is hypothesis testing, the second is estimating conditional distributions.
5.1 Hypothesis Testing
As a first application, we obtain statistical tests of the hypothesis that the true equals the prior density ratio,
[…]
Our third main result provides testable implications of KDM. [^9] Note that (38) implies that $g_{0}=p_{\star}$. However, the converse does not hold, because the orthogonal projection of $g_{\star}-p_{\star}$ onto $\overline{\operatorname{Im}J_{\mathbb{P}}}$ can be zero also when (38) does not hold.
Theorem 5.1.
Assume that the kernel $k$ is universal in the sense that $\overline{\operatorname{Im}J_{\mathbb{P}}}=L^{2}_{\mathbb{P}}$. Then hypothesis (38) holds if and only if
$h_{\lambda}=0,\quad\text{for any $\lambda>0$.}$ (39)
In this case, the following testable properties hold,
  1. (i)
    For any $\eta\in(0,1)$, with sampling probability of at least $1-\eta$, we have
    $\|\big{(}\bm{L}_{\mathbb{P}}^{\top}\bm{L}_{\mathbb{P}}+n\lambda\big{)}^{-1}\big{(}\bm{L}_{\mathbb{Q}}^{\top}\bm{1}-\bm{L}_{\mathbb{P}}^{\top}\bm{p}_{\star}\big{)}\|_{2}\leq\big{(}C_{FS}(\eta,0)+C_{AE}(\epsilon,\lambda)\big{)}\lambda^{-1}n^{-1/2},$ (40)
    for any $\lambda>0$ and $n$, for the coefficients $C_{FS}(\eta,s)$ and $C_{AE}(\epsilon,\lambda)$ given in (20) and (33).
  2. (ii)
    Asymptotically for large $n$, the ${\mathbb{R}}^{m}$-valued sample variable
    $v_{\lambda}\coloneqq n^{-1/2}(\bm{L}_{\mathbb{Q}}^{\top}\bm{1}-\bm{L}_{\mathbb{P}}^{\top}\bm{p}_{\star})\sim{\mathcal{N}}(0,\bm{\Sigma})$ (41)
    is normally distributed with mean zero and $m\times m$-covariance matrix given by
    $\bm{\Sigma}=n^{-1}\bm{L}_{\mathbb{Q}}^{\top}\bm{L}_{\mathbb{Q}}-n^{-2}\bm{L}_{\mathbb{Q}}^{\top}\bm{1}\bm{1}^{\top}\bm{L}_{\mathbb{Q}}+n^{-1}\bm{L}_{\mathbb{P}}^{\top}\operatorname{diag}(\bm{p}_{\star})^{2}\bm{L}_{\mathbb{P}}-n^{-2}\bm{L}_{\mathbb{P}}^{\top}\bm{1}\bm{1}^{\top}\bm{L}_{\mathbb{P}}.$ (42)
  3. (iii)
    Denote by $\bm{\Sigma}=\bm{A}\bm{W}\bm{A}^{\top}$ the spectral decomposition of the matrix in (42), with normalized eigenvectors $\bm{A}=[a_{1},\dots,a_{m}]$ and eigenvalues $w_{1}\geq\cdots\geq w_{m}\geq 0$ on the diagonal of $\bm{W}$. Then, asymptotically for large $n$, for any $\ell\leq m$ such that $w_{\ell}>0$, the test statistic
    $T_{\ell}\coloneqq\sum_{i=1}^{\ell}w_{i}^{-1}\big{(}a_{i}^{\top}v_{\lambda})^{2}\sim\chi^{2}(\ell)$ (43)
    is $\chi^{2}$-distributed with $\ell$ degrees of freedom. [^10] Examples of sample-based choices of $\ell$ include: relative thresholding $\ell=\max\{i:w_{i}\geq tw_{1}\}$, or explained variation $\ell=\min\{i:\sum_{j=1}^{i}w_{j}\geq t\sum_{j=1}^{m}w_{j}\}$, for some $t\in(0,1)$. Other choices are possible and subject to future research.
For $p_{\star}=1$, the hypothesis (38) is equivalent to ${\mathbb{Q}}={\mathbb{P}}$, which is the same hypothesis as studied in Pfister et al. (2017). Then the sample variable $u_{\lambda}$ in (82) is the scaled difference $S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{1}$ between the sample kernel mean embeddings of ${\mathbb{Q}}$ and ${\mathbb{P}}$. As a corollary of Theorem 5.1 we thus recover the kernel two-sample test of Gretton et al. (2012) as a special case, connecting their maximum mean discrepancy approach (Gretton et al., 2012, Theorem 5) with the Hilbert–Schmidt independence criterion (Pfister et al., 2017, Proposition 1).
5.2 Conditional Distribution Estimation
As a second application, we obtain a novel, consistent estimator of conditional distributions. Thereto, we assume that ${\mathcal{Z}}={\mathcal{X}}\times{\mathcal{Y}}$ is the product of two countably generated measurable spaces ${\mathcal{X}}$ and ${\mathcal{Y}}$, and consider random variables $X$ and $Y$ with values in ${\mathcal{X}}$ and ${\mathcal{Y}}$. We denote by ${\mathbb{P}}_{X}$, ${\mathbb{P}}_{Y}$ and ${\mathbb{P}}_{(X,Y)}$ their marginal and joint distributions, and set ${\mathbb{P}}\coloneqq{\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}$ and ${\mathbb{Q}}\coloneqq{\mathbb{P}}_{(X,Y)}$. Assumption (1) then reads
${\mathbb{P}}_{(X,Y)}\ll{\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}.$ (44)
We first derive our estimator of the conditional distribution based on a joint sample of $(X,Y)$. We then briefly discuss the relevance and restrictiveness of Assumption (44). We also relate our conditional distribution estimator to the literature.
5.2.1 Derivation of the Estimator
Assumption (44) implies that the conditional distribution of $Y$ given $X=x$, and of $X$ given $Y=y$, exists and (a version of it) can be expressed in terms of $g_{\star}$ as
${\mathbb{P}}_{Y\mid X=x}(dy)=g_{\star}(x,y){\mathbb{P}}_{Y}(dy),$ (45)
and ${\mathbb{P}}_{X\mid Y=y}(dx)=g_{\star}(x,y){\mathbb{P}}_{X}(dx)$, respectively. In the following, we will concentrate on the former. By symmetry, all statements also apply to the latter.
For any hypothesis density ratio (4) with $h\in{\mathcal{H}}$, we thus obtain the corresponding model of the true conditional expectation ${\mathbb{E}}[f(Y)\mid X=x]=\int_{\mathcal{Y}}f(y){\mathbb{P}}_{Y\mid X=x}(dy)$ given by
${\mathbb{E}}_{h}[f(Y)\mid X=x]\coloneqq\int_{\mathcal{Y}}f(y)(p_{\star}(x,y)+h(x,y))\,{\mathbb{P}}_{Y}(dy),\quad\text{for $f\in L^{2}_{{\mathbb{P}}_{Y}}$.}$ (46)
In view of the isometric isomorphisms
$L^{2}_{{\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}}=L^{2}_{{\mathbb{P}}_{X}}\otimes L^{2}_{{\mathbb{P}}_{Y}}=L^{2}_{{\mathbb{P}}_{X}}({\mathcal{X}};L^{2}_{{\mathbb{P}}_{Y}})=L^{2}_{{\mathbb{P}}_{Y}}({\mathcal{Y}};L^{2}_{{\mathbb{P}}_{X}}),$ (47)
the right hand side of (46) as a function of $x$ is an element in $L^{2}_{{\mathbb{P}}_{X}}$. In that sense, we can decompose the error function (6) as
${\mathcal{E}}(h)^{2}$ $=\int_{\mathcal{X}}\int_{\mathcal{Y}}\big{(}g_{\star}(x,y)-p_{\star}(x,y)-h(x,y)\big{)}^{2}\,{\mathbb{P}}_{Y}(dy){\mathbb{P}}_{X}(dx)$ (48)
$=\int_{\mathcal{X}}\bigg{(}\sup_{\|f\|_{L^{2}_{{\mathbb{P}}_{Y}}}\leq 1}\Big{|}{\mathbb{E}}[f(Y)\mid X=x]-{\mathbb{E}}_{h}[f(Y)\mid X=x]\Big{|}\bigg{)}^{2}\,{\mathbb{P}}_{X}(dx).$
Hence ${\mathcal{E}}(h)$ can be interpreted as the root mean square worst-case error of the conditional expectation model (46). The guarantees in Theorem 4.5 apply accordingly to the following low-rank estimator of the conditional expectation, [^11] In the sense that for any test function $f\in L^{2}_{{\mathbb{P}}_{Y}}$ the $L^{2}_{{\mathbb{P}}_{X}}$-error of the conditional expectation is bounded by $\big{\|}{\mathbb{E}}[f(Y)\mid X=\cdot]-{\mathbb{E}}_{\widetilde{h}_{\lambda}}[f(Y)\mid X=\cdot]\big{\|}_{L^{2}_{{\mathbb{P}}_{X}}}\leq{\mathcal{E}}(\widetilde{h}_{\lambda})\|f\|_{L^{2}_{{\mathbb{P}}_{Y}}}$, which can be bounded by the right hand sides of (35) and (37), respectively. which we obtain by combining (46) with the previous sections,
${\mathbb{E}}_{\widetilde{h}_{\lambda}}[f(Y)\mid X=x]$ $=\int_{\mathcal{Y}}f(y)\big{(}p_{\star}(x,y)+\widetilde{h}_{\lambda}(x,y)\big{)}\,{\mathbb{P}}_{Y}(dy)$ (49)
$\approx\bar{n}^{-1}\sum_{i=1}^{\bar{n}}f(\bar{y}_{i})\big{(}p_{\star}(x,\bar{y}_{i})+\widetilde{h}_{\lambda}(x,\bar{y}_{i})\big{)}$ (50)
$\approx\frac{\sum_{i=1}^{\bar{n}}f(\bar{y}_{i})\big{(}p_{\star}(x,\bar{y}_{i})+\widetilde{h}_{\lambda}(x,\bar{y}_{i})\big{)}^{+}}{\sum_{j=1}^{\bar{n}}\big{(}p_{\star}(x,\bar{y}_{j})+\widetilde{h}_{\lambda}(x,\bar{y}_{j})\big{)}^{+}},$ (51)
for any $f\in L^{2}_{{\mathbb{P}}_{Y}}$. Here we estimate the ${\mathbb{P}}_{Y}$-integral in (49) by the empirical counterpart given by (50), for an auxiliary i.i.d. sample $\bar{y}_{1},\dots,\bar{y}_{\bar{n}}$ of ${\mathbb{P}}_{Y}$. The last approximation in (51) is practically motivated such that we integrate $f$ with respect to a bona-fide conditional distribution. [^12] The density ratio $g_{\star}$ satisfies the implicit structural properties $g_{\star}(x,y)\geq 0$ ${\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}$-a.s., $\int_{\mathcal{Y}}g_{\star}(x,y){\mathbb{P}}_{Y}(dy)=1$ ${\mathbb{P}}_{X}$-a.s., and $\int_{\mathcal{X}}g_{\star}(x,y){\mathbb{P}}_{X}(dx)=1$ ${\mathbb{P}}_{Y}$-a.s.
The estimator $\widetilde{h}_{\lambda}$ in (49) requires i.i.d. samples $z_{{\mathbb{P}},1},\dots,z_{{\mathbb{P}},n}$ of ${\mathbb{P}}\coloneqq{\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}$ and $z_{{\mathbb{Q}},1},\dots,z_{{\mathbb{Q}},n}$ of ${\mathbb{Q}}\coloneqq{\mathbb{P}}_{(X,Y)}$. If we could only sample from the joint distribution, we would consider an i.i.d. sample $(x_{1},y_{1}),\dots,(x_{3n},y_{3n})$ of ${\mathbb{P}}_{(X,Y)}$ of size $3n$ and set
$z_{{\mathbb{P}},i}\coloneqq(x_{2i-1},y_{2i}),\quad z_{{\mathbb{Q}},i}\coloneqq(x_{2n+i},y_{2n+i}),\quad i=1,\dots,n.$ (52)
Remark 5.2.
The sampling scheme (52) may be prohibitive in terms of the required total sample size $3n$. In practice, observing an i.i.d. sample $(x_{1},y_{1}),\dots,(x_{n},y_{n})$ of ${\mathbb{P}}_{(X,Y)}$ of size $n$, one could replace (52) by $z_{{\mathbb{P}},i}\coloneqq(x_{i},y_{i+1})$, with $y_{n+1}\coloneqq y_{1}$, and $z_{{\mathbb{Q}},i}\coloneqq(x_{i},y_{i})$, $i=1,\dots,n$. This comes at the cost of introducing bias though the mutual dependence of $z_{{\mathbb{P}},i}$, $z_{{\mathbb{P}},i+1}$, $z_{{\mathbb{Q}},i}$. Similarly, the auxiliary i.i.d. sample $\bar{y}_{1},\dots,\bar{y}_{\bar{n}}$ of ${\mathbb{P}}_{Y}$ in (50) and (51) could be obtained by setting $\bar{y}_{i}=y_{i}$ for some $\bar{n}\leq n$.
Under appropriate technical conditions, and with similar arguments as below Theorem 4.5, the approximation (51) of (50) can be shown to be asymptotically exact for $\epsilon=\epsilon_{n}\to 0$, $\lambda=\lambda_{n}\to 0$ as $n\to\infty$. This follows from (13) and (34), and because the ${\mathcal{H}}$-norm dominates the sup-norm by assumption (3).
5.2.2 Relevance and Restrictiveness of Assumption (44)
[…]
Lemma 5.3.
Assume there exist measures $\mu_{\mathcal{X}}$ and $\mu_{\mathcal{Y}}$ on ${\mathcal{X}}$ and ${\mathcal{Y}}$ such that ${\mathbb{P}}_{(X,Y)}$ is absolutely continuous with respect to the product measure $\mu_{\mathcal{X}}\otimes\mu_{\mathcal{Y}}$, with say density $f$,
$d{\mathbb{P}}_{(X,Y)}=fd\,(\mu_{\mathcal{X}}\otimes\mu_{\mathcal{Y}}).$ (53)
Then (44) holds.
A simple example where the assumption of Lemma 5.3 holds is easily constructed. Take ${\mathcal{X}}={\mathcal{Y}}={\mathbb{R}}$ and $\mu_{\mathcal{X}}=\mu_{\mathcal{Y}}$ is the Lebesgue measure on ${\mathbb{R}}$, and consider $(X,Y)$ jointly Gaussian with non-degenerate covariance. Then (44) is satisfied. A counter example to (44) is similarly easily devised by setting $Y=X$ for a Gaussian random variable $X$. Then the joint distribution of $(X,Y)=(X,X)$ is concentrated on the diagonal in ${\mathbb{R}}^{2}$ and (44) does not hold. This counter example also illustrates that exceptions to (44) are rather degenerate limiting cases.
5.2.3 Comparison with Related Literature
We relate our conditional distribution estimator (49) to other approaches in the literature.
Kernel conditional mean embeddings
In view of the decomposition (48) it is natural to compare conditional expectations (46) calculated from a hypothesis density ratio (4) to the conditional mean embedding introduced in Song et al. (2009). To facilitate this comparison, we assume in this subsection that ${\mathcal{H}}={\mathcal{H}}_{\mathcal{X}}\otimes{\mathcal{H}}_{\mathcal{Y}}$ is the tensor product of two separable RKHS ${\mathcal{H}}_{\mathcal{X}}$ and ${\mathcal{H}}_{\mathcal{Y}}$ with measurable kernels $k_{\mathcal{X}}$ and $k_{\mathcal{Y}}$ on ${\mathcal{X}}$ and ${\mathcal{Y}}$, respectively. In line with (3), we also assume that $\sup_{x\in{\mathcal{X}}}k_{\mathcal{X}}(x,x)<\infty$ and $\sup_{y\in{\mathcal{Y}}}k_{\mathcal{Y}}(y,y)<\infty$, so that the canonical embeddings $J_{{\mathbb{P}}_{X}}\colon{\mathcal{H}}_{\mathcal{X}}\to L^{2}_{{\mathbb{P}}_{X}}$, $J_{{\mathbb{P}}_{Y}}\colon{\mathcal{H}}_{\mathcal{Y}}\to L^{2}_{{\mathbb{P}}_{Y}}$ are Hilbert–Schmidt operators. We then have $k\big{(}(x_{1},y_{1}),(x_{2},y_{2})\big{)}=k_{\mathcal{X}}(x_{1},x_{2})k_{\mathcal{Y}}(y_{1},y_{2})$, and in view of (47) we can identify the operators $J_{{\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}}=J_{{\mathbb{P}}_{X}}\otimes J_{{\mathbb{P}}_{Y}}$.
Combining this with (2), we obtain kernel embeddings of the conditional distributions (45) in the sense that
$\int_{\mathcal{Y}}f(y){\mathbb{P}}_{Y\mid X=x}(dy)=\langle J_{{\mathbb{P}}_{Y}}f,g_{\star}(x,\cdot)\rangle_{L^{2}_{{\mathbb{P}}_{Y}}}=\langle f,J_{{\mathbb{P}}_{Y}}^{\ast}g_{\star}(x,\cdot)\rangle_{{\mathcal{H}}_{Y}},\quad\text{for $f\in{\mathcal{H}}_{Y}$,}$ (54)
and similarly for ${\mathbb{P}}_{X\mid Y=y}(dx)$. We thus identify from (54) the element $J_{{\mathbb{P}}_{Y}}^{\ast}g_{\star}(x,\cdot)\in{\mathcal{H}}_{Y}$ as the conditional mean embedding $\mu_{Y\mid X=x}$ introduced in Song et al. (2009). More specifically, Song et al. (2009) realize $\mu_{Y\mid X=x}$ through unbounded linear operations between the RKHS ${\mathcal{H}}_{X}$ and ${\mathcal{H}}_{Y}$. A rigorous theory is given by Klebanov et al. (2020), which reveals that, albeit mathematically elegant, the linear operator approach of Song et al. (2009) comes with some practical limitations. First, it requires an elaborate analysis based on sophisticated knife-edge technical assumptions. Second, the inverse problem of recovering ${\mathbb{P}}_{Y\mid X=x}$ from $\mu_{Y\mid X=x}$ is ill-posed, as discussed in Schuster et al. (2020), and left open as a “fruitful avenue of research” in Klebanov et al. (2020). Third, the conditional mean embedding acts through (54) only on functions $f\in{\mathcal{H}}_{\mathcal{Y}}$.
Our approach, via the density ratio $g_{\star}$, overcomes the above limitations of the traditional conditional mean embedding approach. First, our construction of $\mu_{Y\mid X=x}$ is more elementary and feasible under verifiable technical assumptions, such as (1) and (2). Second, our hypothesis model (4) of $g_{\star}$ yields the candidate conditional distribution ${\mathbb{P}}_{Y\mid X=x}$ directly using identity (45). Third, we thus compute conditional expectations (46) directly for all test functions $f\in L^{2}_{{\mathbb{P}}_{Y}}$, and not only for $f\in{\mathcal{H}}_{\mathcal{Y}}$.
Other approaches
Sugiyama et al. (2010); Hinder et al. (2021) estimate conditional density functions with respect to Lebesgue measure. Our setup is more general. In turn, we do not obtain a direct estimator for the conditional Lebesgue density function, hence the evaluation metric in Sugiyama et al. (2010); Hinder et al. (2021) (the negative log likelihood) cannot be applied here and a direct comparison with the scores in Sugiyama et al. (2010); Hinder et al. (2021) is not possible.
Also the setup in Suzuki et al. (2009); Sugiyama (2013) is restricted to continuous variables with a Lebesgue density. In fact, they assume the assumptions of Lemma 5.3 hold for the Lebesgue measures $\mu_{\mathcal{X}}(dx)=dx$ and $\mu_{\mathcal{Y}}(dy)=dy$. In this case, the “density ratio $w(x,y)$” in Suzuki et al. (2009) is equal to our density ratio $g_{\star}(x,y)$. Their estimator of the density ratio is based on a sampling scheme, similar to Filipović et al. (2024), which is different from ours. As a consequence, they do not provide finite sample guarantees nor a central limit theorem.
There is a vast literature on (semi-)parametric distributional regression approaches, which includes structured additive models, see Rügamer et al. (2023) for a recent overview. In contrast, our approach is nonparametric functional analytic and learns the density ratio in a reproducing kernel Hilbert space, which comes with asymptotic theory and finite sample guarantees. Isotonic distribution regression in Henzi et al. (2021) is to date specialized on univariate response variables and conditional distributions, conformal predictive distributions from Vovk et al. (2019) are based on regression problems. Shen and Meinshausen (2024) estimates conditional distributions only.
6 Empirical Experiments
In the following, we develop a number of numerical experiments with both simulated and real data to investigate the properties of KDM across a variety of settings for the applications laid out above. In particular for the real data application, we need to validate the hyperparameters of the RKHS and $\lambda$. To this end, we assume the availability of validation samples $\overline{z}_{{\mathbb{M}},1},\dots,\overline{z}_{{\mathbb{M}},\overline{n}}$ of size $\overline{n}$ of both measures ${\mathbb{M}}={\mathbb{P}},{\mathbb{Q}}$. The validation loss, say ${\mathcal{L}}(\widetilde{h}_{\lambda})$, is obtained by plugging the low-rank estimator $\widetilde{h}_{\lambda}$ given by (31) in the non-regularized objective function in (15), setting $\lambda=0$ and for the training samples replaced by the validation samples. Expressed in coordinates $\bm{\beta}\coloneqq\bm{R}\big{(}\bm{L}_{\mathbb{P}}^{\top}\bm{L}_{\mathbb{P}}+n\lambda\big{)}^{-1}\big{(}\bm{L}_{\mathbb{Q}}^{\top}\bm{1}-\bm{L}_{\mathbb{P}}^{\top}\bm{p}_{\star}\big{)}$ of $\widetilde{h}_{\lambda}$ in (31), the validation loss is given by
${\mathcal{L}}(\widetilde{h}_{\lambda})=-2\big{(}\bm{1}^{\top}k(\overline{\bm{z}}_{\mathbb{Q}},\bm{z}_{\Pi}^{\top})-\bm{p}^{\star\top}k(\overline{\bm{z}}_{\mathbb{P}},\bm{z}_{\Pi}^{\top})\big{)}\bm{\beta}+\bm{\beta}^{\top}k(\bm{z}_{\Pi},\overline{\bm{z}}_{\mathbb{P}}^{\top})k(\overline{\bm{z}}_{\mathbb{P}},\bm{z}_{\Pi}^{\top}))\bm{\beta}$ (55)
using the notation introduced in Section 4.
6.1 Hypothesis Testing
In this section, we perform the statistical test (38) using a battery of statistical distributions taken from Zeng et al. (2018) and Ai et al. (2022). The generation of draws from these distributions is described in Appendix C.
n = 1500
model C HSIC Gauss Laplace Polynomial
IndependentClouds 0.00 0.06 0.06 0.09 0.02
W 1.20 1.00 1.00 1.00 1.00
Diamond 0.70 1.00 1.00 1.00 1.00
Parabola 0.25 1.00 0.98 0.99 1.00
TwoParabola 0.35 1.00 0.99 0.99 1.00
Circle 2.75 1.00 1.00 1.00 1.00
Variance 1.20 1.00 1.00 0.99 0.84
Log 0.18 1.00 1.00 1.00 0.53
n = 3000
model C HSIC Gauss Laplace Polynomial
IndependentClouds 0.00 0.06 0.04 0.05 0.04
W 1.20 1.00 1.00 1.00 1.00
Diamond 0.70 1.00 0.97 0.99 1.00
Parabola 0.25 1.00 0.95 0.98 1.00
TwoParabola 0.35 1.00 0.97 0.97 1.00
Circle 2.75 1.00 1.00 1.00 1.00
Variance 1.20 1.00 1.00 0.99 0.99
Log 0.18 1.00 1.00 1.00 0.79
n = 6000
model C HSIC Gauss Laplace Polynomial
IndependentClouds 0.00 0.06 0.03 0.04 0.03
W 1.20 1.00 1.00 0.99 1.00
Diamond 0.70 1.00 1.00 0.98 1.00
Parabola 0.25 1.00 1.00 0.96 1.00
TwoParabola 0.35 1.00 1.00 0.94 1.00
Circle 2.75 1.00 1.00 0.99 1.00
Variance 1.20 1.00 1.00 0.98 1.00
Log 0.18 1.00 1.00 0.99 0.94
Table 1: Independence testing. The table shows the number of false rejections in the top line (IndependentClouds) at the level of 5%, and the number of correct rejections at the level of 5% for the remaining, dependent, samples, described in Appendix C. The column denoted HSIC shows the results from the Pfister et al. (2017) test using the “gamma” approximation. All models were validated using loss function (55). The numbers represent the empirical probability calculated from $2000$ simulated data sets of length $n=1500,3000,6000$, respectively. Test statistic (43) was truncated at $\ell=\max\{i:w_{i}\geq 10^{-9}w_{1}\}$.
We perform the test using the Gaussian and Laplace kernels (see Rasmussen and Williams, 2005), and polynomial kernel of order $q$ (see Schölkopf and Smola, 2018),
$\quad k_{gauss}(z,z^{\prime})$ $\coloneqq{\rm e}^{-\frac{\|z-z^{\prime}\|^{2}_{2}}{2\,\rho_{gauss}}},$
$\quad k_{laplace}(z,z^{\prime})$ $\coloneqq{\rm e}^{-\rho_{laplace}\|z-z^{\prime}\|_{2}},$
$k_{poly}(z,z^{\prime})$ $\coloneqq(\langle z,z^{\prime}\rangle_{2}+c)^{q}.$
While the Gaussian and Laplacian kernels pertain to infinite-dimensional spaces of smooth, rapidly decaying functions, the polynomial kernel is finite dimensional. All three kernels carry hyperparameters $\rho_{gauss}>0$, $\rho_{laplace}\geq 0$, and $c\geq 0$. Among the three, the polynomial kernel is the only unbounded one, not satisfying (3), and the only one that is not universal. We include it nevertheless in our investigation to test the robustness of the inference results in the previous sections to violations of the assumptions.
Table 1 shows the rejection rates of the test statistic (43) at significance level 5% estimated over $2000$ data sets each, for training sample sizes $n=1500,3000,6000$. The test statistic itself is parameterized by $\ell\leq m$, which determines the number of nonzero eigenvalues of the covariance matrix of the sample variable (41). Here we use the relative thresholding $\ell=\max\{i:w_{i}\geq 10^{-9}w_{1}\}$. All three kernels exhibit better results for larger sample sizes, which is in line with the asymptotic validity of (43). Among the three kernels considered, the Gauss kernel performs best, comparable to the specialized Pfister et al. (2017) HSIC test. While all three kernels show a slight tendency to under-reject, the deviation from the theoretically correct average $p$-value of $5\%$ is quantitatively small. Summarizing, the empirical outcomes of KDM independence testing are encouraging.
6.2 Conditional Distribution Estimation
In this section, we use KDM to estimate conditional distributions in two examples. First, we consider simulated samples generated from mixture distributions. Second, we estimate the conditional distribution of stock returns.
6.2.1 Simulation Study from Mixture Models
We assess the conditional distributions and conditional expectations (51) through scoring rules as proposed by Gneiting and Raftery (2007), and confront KDM with the locally smoothed kernel density estimator (Racine, 2008).
To this end, we generate mixtures from $j=1,2,3$ Gaussian unit variance distributions as follows. Seeding each simulation $i=1,\ldots,200$, we draw random $4\times 4$ correlation matrices using the sampler from Archakov and Hansen (2021). Subsequently, we draw random means $\mu_{1}^{(i,j)}\sim{\mathcal{U}}(-0.2,0.2),\ldots,\mu_{4}^{(i,j)}\sim{\mathcal{U}}(-0.2,0.2)$. These give mixture distributions ${\mathcal{N}}^{(i,j)}$ describing the joint realizations on ${\mathbb{R}}^{2}\times{\mathbb{R}}^{2}$ as follows.
For each simulation $i$, we additionally sample mixture weights $w_{1}^{(i)},\ldots,w_{j}^{(i)}$ from the probability simplex. These weights are then used to generate an i.i.d. sample of size $3n$ as
$(X_{t}^{(i)},Y_{t}^{(i)})\sim{\mathcal{N}}^{(i,J_{t}^{(i)})},\text{ where }J_{t}^{(i)}\coloneqq\operatorname*{arg\,max\,}_{j}\sum_{l=1}^{j}{w_{l}^{(i)}}\leq\epsilon_{t}^{(i)},$
and $\epsilon_{t}^{(i)}\sim{\mathcal{U}}(0,1)$, for $t=1,\dots,3n$. For the $i$-th simulation run, we train KDM with $z_{{\mathbb{Q}},1}^{(i)}=(X_{1}^{(i)},Y_{1}^{(i)}),\dots,z_{{\mathbb{Q}},n}=(X_{n}^{(i)},Y_{n}^{(i)})$, and $z_{{\mathbb{P}},1}=(X_{n+1}^{(i)},Y_{2n+1}^{(i)}),\dots,z_{{\mathbb{P}},n}=(X_{2n}^{(i)},Y_{3n}^{(i)})$, through $20$-fold cross validation with loss function (55).
To benchmark KDM, we estimate a nonparametric locally smoothed kernel density as described in Li and Racine (2006) on the $3n$ data points using the R package np. The package finds the optimal bandwidth that suits the data of simulation run $i$.
The benchmark metric is based on the energy scoring rule, as implemented in the R package scoringRules, for out-of-sample data $\bm{x}_{1},\ldots,\bm{x}_{m}\in{\mathbb{R}}^{2}$ and $\bm{y}\in{\mathbb{R}}^{2}$ generated from the mixture distributions described above,
$ES_{{\mathcal{M}}}(\bm{y})\coloneqq\frac{1}{m}\sum_{i=1}^{m}w^{{\mathcal{M}}}(\bm{x}_{i},\bm{y})\|\bm{y}-\bm{x}_{i}\|_{2}-\frac{1}{2m^{2}}\sum_{i,j=1}^{m}w^{{\mathcal{M}}}(\bm{x}_{i},\bm{y})w^{{\mathcal{M}}}(\bm{x}_{j},\bm{y})\|\bm{x}_{i}-\bm{x}_{j}\|_{2},$
where the weights $w^{{\mathcal{M}}}(\bm{x},\bm{y})$ are taken to be the conditional densities of $\bm{y}|\bm{x}$ evaluated for ${\mathcal{M}}\in\{KDM,np\}$. We then consider the average
energy score $\text{energy score differential}\coloneqq\frac{1}{m}\sum_{i=1}^{m}(ES_{np}(\bm{y}_{i})-ES_{KDM}(\bm{y}_{i})),$
for $\bm{y}_{1},\ldots,\bm{y}_{m}$ from the joint sample $(\bm{x}_{1},\bm{y}_{1}),\ldots,(\bm{x}_{m},\bm{y}_{m})$.
Figure 1 shows the distribution of the energy score differential over simulation runs $i=1,\ldots,200$, and for one, two, and three clusters in the Gaussian mixture for $m=n=1000$. KDM uses the Gaussian kernel for this exercise. The mean energy score differential can be seen to be in favor of KDM. The empirical distributions are also pronouncedly skewed in favor of KDM. Note that the computational effort for bandwidth selection required by np, in particular in larger data sets and higher dimensions is substantial. We therefore refrain from considering such larger data sets.
Figure 1: Energy score differential. The figure shows the distribution of the difference between the energy score computed from KDM and the nonparametric kernel density estimation as described in Li and Racine (2006). The data are generated from a mixture of normal distributions with one, two, or three clusters with $n=1000$. Higher is better.
6.2.2 Conditional Distribution of Stock Returns
To showcase KDM with real and higher-dimensional data, we consider the joint distribution of eleven realized monthly stock portfolio returns $\bm{Y}_{t+1}$, of which five are from Fama and French (2015), four from Hou et al. (2014), one from He et al. (2017), and one is a momentum factor return, along with eleven predictive variables $\bm{X}_{t}$: book-to-market (BM), net-equity-expansion (ntis), inflation growth (infl), stock variance (svar), dividend yield (DP), default yield (DFY), term spread (TMS), civilian unemployment rate (UNRATE), consumption growth (CONSGR), and the Chicago Fed National Activity Index (CFNAI) from Jan 1963 to Dec 2022 (720 months). Both, the returns, and the conditioning covariates, are indexed by time. From these data, we have ${\mathcal{Z}}={\mathbb{R}}^{11}\times{\mathbb{R}}^{11}$, with total dimension adding up to 22. This is too high-dimensional for the locally smoothed nonparametric kernel density estimation from Li and Racine (2006), and we use instead a Gaussian distribution on ${\mathcal{Z}}$ as a benchmark, whose moments are estimated from sample averages, from which we compute conditional moments. We use monthly expanding training windows starting in Jan 1963, with lengths ranging from 200 to 719 months, each followed by one test month (the first test month is Sep 1979, the last is Dec 2022). For KDM, kernel parameters are selected via $k=8$-fold cross-validation using the loss function in (55).
We consider several benchmarks,
$R^{2}_{t,T,\text{OOS}}\coloneqq 1-\frac{\sum_{s=t}^{T-1}\|\bm{Y}_{s+1}-\bm{\mu}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{KDM}}\|_{2}^{2}}{\sum_{s=t}^{T-1}\|\bm{Y}_{s+1}-\bm{\mu}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{{\mathcal{M}}}\|_{2}^{2}},$ (56)
[…]
$R^{2,2}_{t,T,\text{OOS}}\coloneqq 1-\frac{\sum_{s=t}^{T-1}\|\bm{Y}_{s+1}\bm{Y}_{s+1}^{\top}-(\bm{\Sigma}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{KDM}}+\bm{\mu}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{KDM}}(\bm{\mu}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{KDM}})^{\top})\|_{F}^{2}}{\sum_{s=t}^{T-1}\|\bm{Y}_{s+1}\bm{Y}_{s+1}^{\top}-(\bm{\Sigma}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{Gauss}}+\bm{\mu}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{Gauss}}(\bm{\mu}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{Gauss}})^{\top})\|_{F}^{2}},$ (57)
where $\bm{\Sigma}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{KDM}}$ and $\bm{\Sigma}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{Gauss}}$ are conditional covariance matrices computed from KDM and a conditional Gaussian distribution estimated from sample averages, respectively, and $\|\cdot\|_{F}$ denotes the Frobenius norm. We furthermore use the statistical scoring loss $\mathcal{S}:\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}\to\mathbb{R}$ proposed by Dawid and Sebastiani (1999),
$\mathcal{S}(\bm{x},\bm{\mu},\bm{\Sigma})\coloneqq\log\det\bm{\Sigma}+(\bm{x}-\bm{\mu})^{\top}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu}),$
in particular the excess scoring loss
${\mathcal{S}}_{t,T,\text{OOS}}\coloneqq\frac{1}{T-t}\sum_{s=t}^{T-1}\big{(}{\mathcal{S}}(\bm{Y}_{s+1},\bm{\mu}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{Gauss}},\bm{\Sigma}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{Gauss}})-{\mathcal{S}}(\bm{Y}_{s+1},\bm{\mu}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{KDM}},\bm{\Sigma}_{\bm{Y}_{s+1}|\bm{X}_{s}}^{\text{KDM}})\big{)}.$ (58)
Figure 2 shows the expanding out-of-sample $R^{2}_{t,T,\text{OOS}}$ and $R^{2,2}_{t,T,\text{OOS}}$ over time $T$ ranging from first test month ($t+1$ = Sep 1979) to last (Dec 2022). While KDM exhibits higher predictability against sample averages, it fares particularly well in the 1987 crash as well as the COVID-19 pandemic against the conditional Gaussian distribution for both the polynomial, and the Gaussian kernel. Figure 3 shows the logarithm of the expanding out-of-sample excess scoring rule ${\mathcal{S}}_{t,T,\text{OOS}}$ over time $T$, which is uniformly positive. Summing up, KDM performs well empirically for both simulated and real data.
(a) OOS $R^{2}$ (Gauss kernel)
(b) OOS $R^{2}$ (polynomial kernel)
Figure 2: This figure shows out-of-sample $R^{2}$ from (56) and (57) calculated for the Gaussian kernel in Panel 2(a), and the polynomial kernel in Panel 2(b). Higher is better. The data are eleven monthly US stock returns, and eleven conditioning variables from 1963 up to 2022.
Figure 3: This figure shows the logarithm of the excess scoring rule (58). Higher is better. The data are eleven monthly US stock returns, and eleven conditioning variables from 1963 up to 2022.
7 Conclusion
Kernel density machines (KDM) is a comprehensive data-driven framework for estimating the density ratio (Radon-Nikodym derivative) $\frac{d\mathbb{Q}}{d\mathbb{P}}$ from i.i.d. samples of $\mathbb{P}$ and $\mathbb{Q}$ in a reproducing kernel Hilbert hypothesis space. KDM is computable in particular with large data sets, and allows both finite-sample and asymptotic inference. Accordingly, we illustrate its use within hypothesis testing and conditional distribution estimation. Given the reliance of learning problems on the data-generating probability law and the law induced by a model, KDM can be used in many scenarios ranging from transfer learning problems, and generative models to causal inference.
Appendix A Pivoted Cholesky Decomposition
In this appendix, we provide a self-contained proof of Proposition 4.1 and discuss some pivoting strategies
A.1 Proof of Proposition 4.1
We replace Algorithm 1 by the equivalent formulation given in Algorithm 2, which contains the Schur complement matrices in steps (iii), (iv), (vi) in full form.
Algorithm 2 Pivoted incomplete Cholesky decomposition, full form
Input: kernel matrix $\bm{K}$, tolerance $\epsilon\geq 0$.
Initialize $\bm{S}^{(0)}\coloneqq\bm{K}$, $d^{(0)}\coloneqq\operatorname{diag}\bm{S}^{(0)}$, $\bm{L}^{(0)}\coloneqq[\,]$, $\bm{B}^{(0)}\coloneqq[\,]$, $\Pi^{(0)}\coloneqq\emptyset$, $i\coloneqq 0$.
While $\|d^{(i)}\|_{1}>\epsilon$ (and hence $d^{(i)}\neq\bm{0}$) do the following steps:
  1. (i)
    select a pivot index $\pi_{i+1}\in\{1,\dots,2n\}\setminus\Pi^{(i)}$ such that $d^{(i)}_{\pi_{i+1}}\neq 0$;
  2. (ii)
    set $\Pi^{(i+1)}\coloneqq\Pi^{(i)}\cup\{\pi_{i+1}\}$;
  3. (iii)
    set $\ell_{i+1}\coloneqq(d^{(i)}_{\pi_{i+1}})^{-1/2}\bm{S}^{(i)}e_{\pi_{i+1}}$;
  4. (iv)
    set $b_{i+1}\coloneqq(d^{(i)}_{\pi_{i+1}})^{-1/2}(I_{2n}-\bm{B}^{(i)}{\bm{L}^{(i)}}^{\top})e_{\pi_{i+1}}$;
  5. (v)
    set $\bm{L}^{(i+1)}\coloneqq[\bm{L}^{(i)},\ell_{i+1}]$, $\bm{B}^{(i+1)}\coloneqq[\bm{B}^{(i)},b_{i+1}]$;
  6. (vi)
    for the Schur complement $\bm{S}^{(i+1)}\coloneqq\bm{S}^{(i)}-\ell_{i+1}\ell_{i+1}^{\top}=\bm{K}-\bm{L}^{(i+1)}{\bm{L}^{(i+1)}}^{\top}$ denote its diagonal by $d^{(i+1)}\coloneqq\operatorname{diag}\bm{S}^{(i+1)}=d^{(i)}-\ell_{i+1}\circ\ell_{i+1}$.
  7. (vii)
    set $i\coloneqq i+1$
Output: rank $m\coloneqq i$, pivot set $\Pi\coloneqq\Pi^{(m)}$, matrices $\bm{L}\coloneqq\bm{L}^{(m)}$, $\bm{B}\coloneqq\bm{B}^{(m)}$
The following arguments assert that Algorithm 2 works. First, note that every Schur complement $\bm{S}^{(i+1)}$ in step (vi) is symmetric positive semidefinite with trace given by $\|d^{(i+1)}\|_{1}$, see Harbrecht et al. (2012, Lemma 2.1). We claim that
$\bm{K}\bm{B}^{(i+1)}$ $=\bm{L}^{(i+1)},$ (59)
${\bm{B}^{(i+1)}}^{\top}\bm{L}^{(i+1)}$ $=I_{i+1},$ (60)
$\operatorname{Im}\bm{B}^{(i+1)}$ $=\operatorname{span}\{e_{\pi_{1}},\dots,e_{\pi_{i+1}}\},$ (61)
$d^{(i+1)}_{j}=\bm{S}^{(i+1)}_{jk}=\bm{S}^{(i+1)}_{kj}$ $=0\quad\text{for all $j\in\Pi^{(i+1)}$ and $k\in\{1,\dots,2n\}$.}$ (62)
From (62) it follows that $\ell_{i+1,j}=0$ for all $j\in\Pi^{(i)}$.
We prove properties (59), (60), (61), and (62) by induction. For the base case $i=0$, we have $\bm{L}^{(1)}=\ell_{1}=(\bm{K}_{\pi_{1},\pi_{1}})^{-1/2}\bm{K}_{:,\pi_{1}}$, $\bm{B}^{(1)}=b_{1}=(\bm{K}_{\pi_{1},\pi_{1}})^{-1/2}e_{\pi_{1}}$. Hence (59), (60), (61) hold by inspection, and we obtain $\bm{S}^{(1)}_{\pi_{1},\pi_{1}}=0$, which proves (62).
For the induction step $i\mapsto i+1$, we first calculate
$\bm{K}b_{i+1}$ $=(d^{(i)}_{\pi_{i+1}})^{-1/2}(\bm{K}-\bm{K}\bm{B}^{(i)}{\bm{L}^{(i)}}^{\top})e_{\pi_{i+1}}$
$=(d^{(i)}_{\pi_{i+1}})^{-1/2}(\bm{K}-\bm{L}^{(i)}{\bm{L}^{(i)}}^{\top})e_{\pi_{i+1}}=(d^{(i)}_{\pi_{i+1}})^{-1/2}\bm{S}^{(i)}e_{\pi_{i+1}}=\ell_{i+1}.$
Hence $\bm{K}\bm{B}^{(i+1)}=[\bm{K}\bm{B}^{(i)},\bm{K}b_{i+1}]=[\bm{L}^{(i)},\ell_{i+1}]=\bm{L}^{(i+1)}$, which proves (59).
Next we calculate ${\bm{B}^{(i)}}^{\top}\bm{S}^{(i)}={\bm{B}^{(i)}}^{\top}\bm{K}-{\bm{B}^{(i)}}^{\top}\bm{L}^{(i)}{\bm{L}^{(i)}}^{\top}={\bm{L}^{(i)}}^{\top}-{\bm{L}^{(i)}}^{\top}=\bm{0}$. Hence
$b_{i+1}^{\top}\ell_{i+1}=(d^{(i)}_{\pi_{i+1}})^{-1}e_{\pi_{i+1}}^{\top}(I_{2n}-\bm{L}^{(i)}{\bm{B}^{(i)}}^{\top})\bm{S}^{(i)}e_{\pi_{i+1}}^{\top}=(d^{(i)}_{\pi_{i+1}})^{-1}\bm{S}^{(i)}_{\pi_{i+1},\pi_{i+1}}=1,$
and ${\bm{B}^{(i)}}^{\top}\ell_{i+1}=(d^{(i)}_{\pi_{i+1}})^{-1/2}{\bm{B}^{(i)}}^{\top}\bm{S}^{(i)}e_{\pi_{i+1}}=0$. Moreover,
$b_{i+1}^{\top}\bm{L}^{(i)}=(d^{(i)}_{\pi_{i+1}})^{-1/2}e_{\pi_{i+1}}^{\top}(I_{2n}-\bm{L}^{(i)}{\bm{B}^{(i)}}^{\top})\bm{L}^{(i)}=(d^{(i)}_{\pi_{i+1}})^{-1/2}e_{\pi_{i+1}}^{\top}(\bm{L}^{(i)}-\bm{L}^{(i)})=\bm{0}.$
Combining the above, we derive
${\bm{B}^{(i+1)}}^{\top}\bm{L}^{(i+1)}=\begin{bmatrix}{\bm{B}^{(i)}}^{\top}\bm{L}^{(i)}&{\bm{B}^{(i)}}^{\top}\ell_{i+1}\\ b_{i+1}^{\top}\bm{L}^{(i)}&b_{i+1}^{\top}\ell_{i+1}\end{bmatrix}=\begin{bmatrix}I_{i}&\bm{0}\\ \bm{0}&1\end{bmatrix},$
which proves (60).
Next, we observe that $b_{i+1}=(d^{(i)}_{\pi_{i+1}})^{-1/2}(e_{\pi_{i+1}}-\bm{B}^{(i)}{\bm{L}^{(i)}}^{\top}e_{\pi_{i+1}})$ lies in the span of $\{e_{\pi_{1}},\dots,e_{\pi_{i+1}}\}$. As $\bm{B}^{(i+1)}$ has full rank, by (60), this proves (61). Finally, by induction hypothesis, we have $\ell_{i+1,j}=0$ for all $j\in\Pi^{(i)}$ and $\ell_{i+1,\pi_{i+1}}=(d^{(i)}_{\pi_{i+1}})^{1/2}$. Hence $d^{(i+1)}_{j}=d^{(i)}_{j}-0=0$ for all $j\in\Pi^{(i)}$ and $d^{(i+1)}_{\pi_{i+1}}=d^{(i)}_{\pi_{i+1}}-d^{(i)}_{\pi_{i+1}}=0$, which proves (62).
It remains to observe that Algorithm 2 can be implemented without computing the full Schur complement matrices $\bm{S}^{(i+1)}$. In every iteration step, we only need to compute and store $\bm{L}^{(i+1)}$, $\bm{B}^{(i+1)}$, and $d^{(i+1)}$. In fact, we can replace steps (iii), (iv), (vi) by the corresponding steps (iii), (iv), (vi) in Algorithm 1.
Property (61) now reveals that only $m$ rows of $\bm{B}$ are different from zero, and they are captured by the $m\times m$-matrix $\bm{R}$. Properties (22)–(23) then follow directly from (59)–(61). Similarly, it follows that the column vectors $\bm{K}_{:,\Pi}$ are linearly independent and (25)–(26) hold. The trace error bound (24) holds by the stopping criterion of the algorithm. This completes the proof of Proposition 4.1.
Remark A.1.
As a corollary of Proposition 4.1, for $\epsilon=0$, we obtain the complete Cholesky decomposition $\bm{K}=\bm{L}\bm{L}^{\top}$ for $m=\operatorname{rank}(\bm{K})$. [^13] Modulo a permutation of the columns of $\bm{L}$ such that it becomes lower triangular. As a consequence, the Schur complement vanishes $\bm{S}^{(i+1)}=\bm{0}$ for $i+1\geq\operatorname{rank}(\bm{K})$ in Algorithm 2.
A.2 Pivoting Strategies
There exist many pivoting strategies for step (i). An optimal greedy choice would be to maximally reduce the trace of the Schur complement $\bm{S}^{(i+1)}$, which according to step (vi) is given by $\operatorname{trace}(\bm{S}^{(i+1)})=\|d^{(i)}\|_{1}-\|\ell_{i+1}\|_{2}^{2}$. This is tantamount to maximize $\|\ell_{i+1}\|_{2}^{2}$. However, this would require the knowledge of the Euclidean norms of all columns on $\bm{S}^{(i)}$ and thus the evaluation of the full matrix $\bm{K}$, which may be computationally prohibitive. We opt here for a greedy version where we simply remove the maximal element on the diagonal $d^{(i)}$ of $\bm{S}^{(i)}$,
$\pi_{i+1}=\operatorname*{arg\,max\,}_{j}d^{(i)}_{j}.$ (63)
Alternative choices, such as random pivoting, are discussed in Chen et al. (2023, Section 2.3). Another alternative, which falls outside the scope of this paper, is presented in the following subsection.
A.3 Orthogonal Matching Pursuit
An alternative pivoting strategy is orthogonal matching pursuit, see, e.g., Vincent and Bengio (2002). Here we assume a target function $f\in{\mathcal{H}}$ be given. The choice of $\Pi$ is such that the subspace ${\mathcal{H}}_{\Pi}$ given in Corollary 4.2 optimally approximates $f$ in the sense that $\|f-P_{\Pi}f\|_{\mathcal{H}}^{2}$ be minimal. Equivalently, such that $\|P_{\Pi}f\|_{\mathcal{H}}^{2}=\|\bm{B}^{\top}f(\bm{z})\|_{2}^{2}=\sum_{i=1}^{m}(b_{i}^{\top}f(\bm{z}))^{2}$ be maximal. A greedy pivoting strategy would choose $\pi_{i+1}$ that maximizes $(b_{i+1}^{\top}f(\bm{z}))^{2}$. We can efficiently compute this as
$b_{i+1}^{\top}f(\bm{z})=(d^{(i)}_{\pi_{i+1}})^{-1/2}e_{\pi_{i+1}}^{\top}\big{(}I_{2n}-\bm{L}^{(i)}{\bm{B}^{(i)}}^{\top}\big{)}f(\bm{z})=(d^{(i)}_{\pi_{i+1}})^{-1/2}\big{(}f(z_{\pi_{i+1}})-w_{i,\pi_{i+1}}\big{)},$ (64)
where we recursively set $w_{i+1}\coloneqq w_{i}+\ell_{i+1}b_{i+1}^{\top}f(\bm{z})$ with $w_{0}\coloneqq 0$. This strategy again has complexity ${\mathcal{O}}(m^{2}n)$. Hence, we set
$\pi_{i+1}\coloneqq\operatorname*{arg\,max\,}_{j:d^{(i)}_{j}\geq\eta_{i+1}}(d^{(i)}_{j})^{-1}\big{(}f(z_{j})-w_{i,j}\big{)}^{2},$ (65)
for some exogenously chosen threshold $\eta_{i+1}>0$. Theoretically one could set $\eta_{i+1}=0$. In practice, however, $\eta_{i+1}$ should be large enough to avoid numerical instability. An example is to let $\eta_{i+1}$ be a specific quantile (e.g., $90\%$) of the distribution of the non-zero diagonal elements $d^{(i)}_{j}\neq 0$. Note that the vector $\big{(}I_{2n}-\bm{L}^{(i)}{\bm{B}^{(i)}}^{\top}\big{)}f(\bm{z})=f^{(i)}(\bm{z})$ in (64) contains just the sample values of the residual $f^{(i)}\coloneqq f-P_{\Pi^{(i)}}f$ of $f$ projected on ${\mathcal{H}}_{\Pi^{(i)}}=\operatorname{span}\{k(\cdot,z_{\pi_{1}}),\dots,k(\cdot,z_{\pi_{i}})\}$ in ${\mathcal{H}}$. Hence we can write $b_{i+1}^{\top}f(\bm{z})=(d^{(i)}_{\pi_{i+1}})^{-1/2}f^{(i)}(z_{\pi_{i+1}})$.
Appendix B Proofs
This appendix collects all remaining proofs from the main text. The proof of Proposition 4.1 is given in Appendix A.
B.1 Proof of Lemma 2.2
For any such $f$ we have $fg_{\star}\in L^{2}_{{\mathbb{P}}}$, $f\in L^{2}_{{\mathbb{Q}}}$, and
$J_{\mathbb{P}}^{\ast}(fg_{\star})=\int_{{\mathcal{Z}}}k(\cdot,z)f(z)g_{\star}(z)\,{\mathbb{P}}(d{z})=\int_{{\mathcal{Z}}}k(\cdot,z)f(z)\,{\mathbb{Q}}(d{z})=J_{\mathbb{Q}}^{\ast}f,$
as desired.
B.2 Proof of Lemma 2.3
(i): This follows from the definition of the projection.
(ii): In view of Lemma 2.2, and by the projection property, we can rewrite (8) as
$h_{\lambda}=(J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}+\lambda)^{-1}J_{\mathbb{P}}^{\ast}(g_{\star}-p_{\star})=(J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}+\lambda)^{-1}J_{\mathbb{P}}^{\ast}(g_{0}-p_{\star}).$ (66)
Convergence now follows as in the proof of Boudabsa and Filipović (2022, Lemma 2.3).
(iii): The first statement is elementary. By assumption we can rewrite (66) as $h_{\lambda}=(J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}+\lambda)^{-1}J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}h_{0}$. Using the notation and setup as in Boudabsa and Filipović (2022, Appendix B.5 and B.6), we can expand $h_{\lambda}$ as $h_{\lambda}=\sum_{i\in I}\frac{\mu_{i}}{\mu_{i}+\lambda}\langle h_{0},u_{i}\rangle_{\mathcal{H}}u_{i}$, and $J_{\mathbb{P}}h_{\lambda}=\sum_{i\in I}\mu_{i}^{1/2}\frac{\mu_{i}}{\mu_{i}+\lambda}\langle h_{0},u_{i}\rangle_{\mathcal{H}}v_{i}$, where $\{u_{i}\}_{i\in I}$ is an orthonormal system in ${\mathcal{H}}$ of eigenfunctions of $J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}$ with eigenvalues $\mu_{i}>0$, for a countable index set $I$. The functions $v_{i}=\mu_{i}^{-1/2}J_{\mathbb{P}}u_{i}$, in turn, form an orthonormal system in $L^{2}_{{\mathbb{P}}}$ of eigenfunctions of $J_{\mathbb{P}}J_{\mathbb{P}}^{\ast}$ with the same eigenvalues $\mu_{i}$. Properties (12) and (13) now follow from dominated convergence, and (14) follows from
$\|J_{\mathbb{P}}h_{0}-J_{\mathbb{P}}h_{\lambda}\|_{L^{2}_{{\mathbb{P}}}}^{2}=\sum_{i\in I}\mu_{i}\frac{\lambda^{2}}{(\mu_{i}+\lambda)^{2}}\langle h_{0},u_{i}\rangle_{\mathcal{H}}^{2}=\lambda\sum_{i\in I}\frac{\mu_{i}\lambda}{(\mu_{i}+\lambda)^{2}}\langle h_{0},u_{i}\rangle_{\mathcal{H}}^{2}\leq\lambda\frac{1}{2}\|h_{0}\|_{\mathcal{H}}^{2},$
where we used the elementary inequality $\mu_{i}\lambda\leq\frac{1}{2}(\mu_{i}+\lambda)^{2}$.
B.3 Proof of Theorem 3.1
The proof of Theorem 3.1 relies on the following key lemma.
Lemma B.1.
  1. (i)
    The ${\mathcal{H}}$-valued sample variable defined by
    ${u_{\lambda}}\coloneqq n^{-1}(S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}(\bm{p}_{\star}+S_{\mathbb{P}}h_{\lambda}))-(J_{\mathbb{Q}}^{\ast}1-J_{\mathbb{P}}^{\ast}(p_{\star}+J_{\mathbb{P}}h_{\lambda}))$ (67)
    has mean zero and ${u_{\lambda}}\to 0$ in ${\mathcal{H}}$ a.s. as $n\to\infty$.
  2. (ii)
    $n^{1/2}{u_{\lambda}}\to{\mathcal{N}}(0,Q_{\lambda})$ in distribution as $n\to\infty$, where the covariance operator $Q_{\lambda}$ is given in (18).
  3. (iii)
    For any $\eta\in(0,1)$, with sampling probability of at least $1-\eta$, we have $\|{u_{\lambda}}\|_{\mathcal{H}}\leq C_{FS}(\eta,\|h_{\lambda}\|_{\mathcal{H}})n^{-1/2}$, for the coefficient $C_{FS}(\eta,s)$ given in (20).
  4. (iv)
    $(n^{-1}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}+\lambda)^{-1}\to(J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}+\lambda)^{-1}$ in operator norm a.s. as $n\to\infty$.
Proof.
(i): We can write ${u_{\lambda}}=n^{-1}\sum_{i=1}^{n}\xi_{i}$, where
$\xi_{i}\coloneqq k(\cdot,z_{{\mathbb{Q}},i})-(p_{\star}(z_{{\mathbb{P}},i})+h_{\lambda}(z_{{\mathbb{P}},i}))k(\cdot,z_{{\mathbb{P}},i})-(J_{\mathbb{Q}}^{\ast}1-J_{\mathbb{P}}^{\ast}(p_{\star}+J_{\mathbb{P}}h_{\lambda}))$ (68)
are i.i.d. ${\mathcal{H}}$-valued random variables with zero mean under the sampling measure, say, $\bm{P}\coloneqq({\mathbb{P}}\otimes{\mathbb{Q}})^{\otimes\infty}$. In view of (3), we obtain the bounds on the operator norms
$\|J_{\mathbb{P}}\|,\,\|J_{\mathbb{P}}^{\ast}\|,\,\|J_{\mathbb{Q}}\|,\,\|J_{\mathbb{Q}}^{\ast}\|\leq\sqrt{\kappa_{\infty}}.$ (69)
Using (5), we obtain that the random variables
$\|\xi_{i}\|_{\mathcal{H}}$ $\leq\|k(\cdot,z_{{\mathbb{Q}},i})\|_{\mathcal{H}}+|p_{\star}(z_{{\mathbb{P}},i})|\|k(\cdot,z_{{\mathbb{P}},i})\|_{\mathcal{H}}+|\langle h_{\lambda},k(\cdot,z_{{\mathbb{P}},i})\rangle_{\mathcal{H}}|\|k(\cdot,z_{{\mathbb{P}},i})\|_{\mathcal{H}}$ (70)
$\quad+\|J_{\mathbb{Q}}^{\ast}\|+\|J_{\mathbb{P}}^{\ast}\|\pi_{\infty}+\|J_{\mathbb{P}}^{\ast}\|\|J_{\mathbb{P}}\|\|h_{\lambda}\|_{\mathcal{H}}$
$\leq 2(\sqrt{\kappa_{\infty}}+\pi_{\infty}\sqrt{\kappa_{\infty}}+\|h_{\lambda}\|_{\mathcal{H}}\kappa_{\infty})\eqqcolon c_{\xi}$
are uniformly bounded. The claim now follows from the law of large numbers, $n^{-1}\sum_{i=1}^{n}\xi_{i}\to 0$ a.s., see Hoffmann-Jorgensen and Pisier (1976, Theorem 2.1).
(ii): Using the above, the functional central limit theorem applies such that $n^{-1/2}\sum_{i=1}^{n}\xi_{i}\to{\mathcal{N}}(0,Q_{\lambda})$ in distribution, see Hoffmann-Jorgensen and Pisier (1976, Theorem 3.6). The covariance operator $Q_{\lambda}$ is given by its action on test functions $f,g\in{\mathcal{H}}$,
$\langle Q_{\lambda}f,g\rangle_{\mathcal{H}}$ $={\mathbb{E}}_{\bm{P}}[\langle\xi_{i},f\rangle_{\mathcal{H}}\langle\xi_{i},g\rangle_{\mathcal{H}}]={\mathbb{E}}_{\bm{P}}[\langle\xi_{{\mathbb{Q}},i},f\rangle_{\mathcal{H}}\langle\xi_{{\mathbb{Q}},i},g\rangle_{\mathcal{H}}]+{\mathbb{E}}_{\bm{P}}[\langle\xi_{{\mathbb{P}},i},f\rangle_{\mathcal{H}}\langle\xi_{{\mathbb{P}},i},g\rangle_{\mathcal{H}}]$
$=\langle J_{\mathbb{Q}}f,J_{\mathbb{Q}}g\rangle_{L^{2}_{{\mathbb{Q}}}}-\langle f,J_{\mathbb{Q}}^{\ast}1\rangle_{\mathcal{H}}\langle g,J_{\mathbb{Q}}^{\ast}1\rangle_{\mathcal{H}}$
$\quad+\big{\langle}(p_{\star}+J_{\mathbb{P}}h_{\lambda})J_{\mathbb{P}}f,(p_{\star}+J_{\mathbb{P}}h_{\lambda})J_{\mathbb{P}}g\big{\rangle}_{L^{2}_{{\mathbb{P}}}}$
$\quad-\langle f,J_{\mathbb{P}}^{\ast}(p_{\star}+J_{\mathbb{P}}h_{\lambda})\rangle_{\mathcal{H}}\langle g,J_{\mathbb{P}}^{\ast}(p_{\star}+J_{\mathbb{P}}h_{\lambda})\rangle_{\mathcal{H}}$
$=\langle J_{\mathbb{Q}}^{\ast}J_{\mathbb{Q}}f,g\rangle_{\mathcal{H}}-\big{\langle}\langle f,J_{\mathbb{Q}}^{\ast}1\rangle_{\mathcal{H}}J_{\mathbb{Q}}^{\ast}1,g\big{\rangle}_{\mathcal{H}}$
$\quad+\big{\langle}J_{\mathbb{P}}^{\ast}\operatorname{diag}(p_{\star}+J_{\mathbb{P}}h_{\lambda})^{2}J_{\mathbb{P}}f,g\rangle_{\mathcal{H}}-\big{\langle}\langle f,J_{\mathbb{P}}^{\ast}(p_{\star}+J_{\mathbb{P}}h_{\lambda})\rangle_{\mathcal{H}}J_{\mathbb{P}}^{\ast}(p_{\star}+J_{\mathbb{P}}h_{\lambda}),g\big{\rangle}_{\mathcal{H}}.$
where we decompose $\xi_{i}=\xi_{{\mathbb{Q}},i}-\xi_{{\mathbb{P}},i}$ into the components $\xi_{{\mathbb{Q}},i}\coloneqq k(\cdot,z_{{\mathbb{Q}},i})-J_{\mathbb{Q}}^{\ast}1$ and $\xi_{{\mathbb{P}},i}\coloneqq(p_{\star}(z_{{\mathbb{P}},i})+h_{\lambda}(z_{{\mathbb{P}},i}))k(\cdot,z_{{\mathbb{P}},i})-J_{\mathbb{P}}^{\ast}(p_{\star}+J_{\mathbb{P}}h_{\lambda})$, which have mean zero and are independent under the sampling measure $\bm{P}$. This proves (18).
(iii): Using the above, Hoeffding’s inequality Pinelis (1994, Theorem 3.5) applies, which bounds the tail probabilities, $\bm{P}\big{[}\|{u_{\lambda}}\|_{\mathcal{H}}>\tau\big{]}\leq 2{\rm e}^{-\frac{\tau^{2}n}{2c_{\xi}^{2}}}$, for any $\tau\geq 0$. This proves the claim.
(iv): This follows as in Boudabsa and Filipović (2022, Lemma B.2). ∎
We can now prove Theorem 3.1. We define $A=J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}+\lambda$ and $b=J_{\mathbb{Q}}^{\ast}1-J_{\mathbb{P}}^{\ast}p_{\star}$ and the sample analogues $\hat{A}=n^{-1}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}+\lambda$ and $\hat{b}=n^{-1}(S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{p}_{\star})$. Using (17) and (10), we then decompose
$\hat{h}_{\lambda}-h_{\lambda}$ $=\hat{A}^{-1}\hat{b}-A^{-1}b=\hat{A}^{-1}(\hat{b}-b)-(A^{-1}-\hat{A}^{-1})b$
$=\hat{A}^{-1}(\hat{b}-b)-\hat{A}^{-1}(\hat{A}-A)A^{-1}b=\hat{A}^{-1}\big{(}\hat{b}-b-(\hat{A}-A)h_{\lambda}\big{)}.$
We obtain
$\hat{h}_{\lambda}-h_{\lambda}=(n^{-1}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}+\lambda)^{-1}{u_{\lambda}}.$ (71)
where ${u_{\lambda}}$ is defined in (67). Part (i) now follows from Lemma B.1(i) and (iv). Part (ii) follows from Lemma B.1(ii) and (iv) and Slutsky’s theorem. Part (iii) follows from Lemma B.1(iii) and using that (71) implies $\|\hat{h}_{\lambda}-h_{\lambda}\|_{\mathcal{H}}\leq\lambda^{-1}\|{u_{\lambda}}\|_{\mathcal{H}}$. This completes the proof of Theorem 3.1.
B.4 Proof of Lemma 4.3
From Proposition 4.1 and Corollary 4.2 we obtain the following matrix representations, for $h\in{\mathcal{H}}$ and $v\in{\mathbb{R}}^{n}$,
$S_{\mathbb{M}}P_{\Pi}h$ $=k(\bm{z}_{\mathbb{M}},\bm{z}_{\Pi}^{\top})\bm{R}\langle\bm{\psi},h\rangle_{\mathcal{H}}=\bm{L}_{\mathbb{M}}\langle\bm{\psi},h\rangle_{\mathcal{H}},$ (72)
$P_{\Pi}S_{\mathbb{M}}^{\ast}v$ $=\bm{\psi}^{\top}\langle\bm{R}^{\top}k(\cdot,\bm{z}_{\Pi}),k(\cdot,\bm{z}_{\mathbb{M}}^{\top})v\rangle_{\mathcal{H}}=\bm{\psi}^{\top}\bm{R}^{\top}\bm{K}_{\Pi,{\mathbb{M}}}v=\bm{\psi}^{\top}\bm{L}_{\mathbb{M}}^{\top}v,$ (73)
$P_{\Pi}S_{\mathbb{M}}^{\ast}S_{\mathbb{M}}P_{\Pi}h$ $=\bm{\psi}^{\top}(\bm{L}_{\mathbb{M}}^{\top}\bm{L}_{\mathbb{M}})\langle\bm{\psi},h\rangle_{\mathcal{H}},$ (74)
$S_{\mathbb{M}}P_{\Pi}S_{\mathbb{M}}^{\ast}v$ $=\bm{L}_{\mathbb{M}}\bm{L}_{\mathbb{M}}^{\top}v.$ (75)
The lemma now follows from (73), (74), and (27).
B.5 Proof of Lemma 4.4
Similarly as for (72)–(75) we derive the matrix representations, for $h\in{\mathcal{H}}$ and $v\in{\mathbb{R}}^{2n}$,
$SP_{\Pi}h$ $=\bm{L}\langle\bm{\psi},h\rangle_{\mathcal{H}},$ (76)
$P_{\Pi}S^{\ast}v$ $=\bm{\psi}^{\top}\bm{L}^{\top}v,$ (77)
$SP_{\Pi}S^{\ast}v$ $=\bm{L}\bm{L}^{\top}v.$ (78)
We define $\bm{q}_{\star}\coloneqq\begin{bmatrix}-\bm{p}_{\star}\\ \bm{1}\end{bmatrix}$, so that we can write $S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{p}_{\star}=S^{\ast}\bm{q}_{\star}$. We decompose
$\hat{h}_{\lambda}-\widetilde{h}_{\lambda}=(S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}+n\lambda)^{-1}S^{\ast}\bm{q}_{\star}-(P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}+n\lambda)^{-1}P_{\Pi}S^{\ast}\bm{q}_{\star}=f_{1}+f_{2},$ (79)
for the two functions
$f_{1}$ $\coloneqq(S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}+n\lambda)^{-1}(S^{\ast}-P_{\Pi}S^{\ast})\bm{q}_{\star},$
$f_{2}$ $\coloneqq\big{(}(S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}+n\lambda)^{-1}-(P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}+n\lambda)^{-1}\big{)}P_{\Pi}S^{\ast}\bm{q}_{\star},$
and derive bounds on $\|f_{1}\|_{\mathcal{H}}$ and $\|f_{2}\|_{\mathcal{H}}$.
First, combining the operator norm identity $\|A\|^{2}=\|A^{\ast}A\|$ and the equality $(S^{\ast}-P_{\Pi}S^{\ast})^{\ast}(S^{\ast}-P_{\Pi}S^{\ast})=SS^{\ast}-SP_{\Pi}S^{\ast}$, and using (78), we obtain the operator norm bound
$\|S^{\ast}-P_{\Pi}S^{\ast}\|^{2}=\|SS^{\ast}-SP_{\Pi}S^{\ast}\|=\|\bm{K}-\bm{L}\bm{L}^{\top}\|\leq\operatorname{trace}(\bm{K}-\bm{L}\bm{L}^{\top})\leq\epsilon.$
On the other hand, we have $\|\bm{q}_{\star}\|_{2}^{2}=\sum_{i=1}^{n}(\bm{p}_{\star,i}^{2}+1)\leq n(\pi_{\infty}^{2}+1)$, and hence
$\|f_{1}\|_{\mathcal{H}}\leq(n\lambda)^{-1}\epsilon^{1/2}\|\bm{q}_{\star}\|_{2}\leq(n\lambda)^{-1}\epsilon^{1/2}n^{1/2}(\pi_{\infty}+1),$ (80)
using the operator norm bound $\|(A^{\ast}A+\lambda)^{-1}\|\leq\lambda^{-1}$. Second, we rearrange terms
$f_{2}$ $=(S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}+n\lambda)^{-1}\big{(}P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}-S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}\big{)}(P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}+n\lambda)^{-1}P_{\Pi}S^{\ast}\bm{q}_{\star}$
$=(S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}+n\lambda)^{-1}\big{(}P_{\Pi}S_{\mathbb{P}}^{\ast}-S_{\mathbb{P}}^{\ast}\big{)}S_{\mathbb{P}}P_{\Pi}(P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}+n\lambda)^{-1}S^{\ast}\bm{q}_{\star}.$
Similarly as above, we derive the bounds on the following operator norms
$\|P_{\Pi}S_{\mathbb{P}}^{\ast}-S_{\mathbb{P}}^{\ast}\|^{2}=\|S_{\mathbb{P}}S_{\mathbb{P}}^{\ast}-S_{\mathbb{P}}P_{\Pi}S_{\mathbb{P}}^{\ast}\|=\|\bm{K}_{\mathbb{P}}-\bm{L}_{\mathbb{P}}\bm{L}_{\mathbb{P}}^{\top}\|\leq\operatorname{trace}(\bm{K}-\bm{L}\bm{L}^{\top})\leq\epsilon,$
using (75), and
$\|S_{\mathbb{P}}P_{\Pi}(P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}+n\lambda)^{-1}\|^{2}=\|(P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}+n\lambda)^{-1}P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}(P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}+n\lambda)^{-1}\|\\ \leq\|(P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}+n\lambda)^{-1}\|\|P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}(P_{\Pi}S_{\mathbb{P}}^{\ast}S_{\mathbb{P}}P_{\Pi}+n\lambda)^{-1}\|\leq(n\lambda)^{-1},$
using the operator norm bound $\|A^{\ast}A(A^{\ast}A+\lambda)^{-1}\|\leq 1$. On the other hand, we have $\|S^{\ast}\bm{q}_{\star}\|_{\mathcal{H}}\leq\sum_{i=1}^{n}(\|k(\cdot,z_{{\mathbb{P}},i})\|_{\mathcal{H}}\bm{p}_{\star,i}+\|k(\cdot,z_{{\mathbb{Q}},i})\|_{\mathcal{H}}\cdot 1)\leq n\kappa_{\infty}^{1/2}(\pi_{\infty}+1)$, and hence
$\|f_{2}\|_{\mathcal{H}}\leq(n\lambda)^{-1}\epsilon^{1/2}(n\lambda)^{-1/2}\|S^{\ast}\bm{q}_{\star}\|_{\mathcal{H}}\leq(n\lambda)^{-1}\epsilon^{1/2}n^{1/2}\lambda^{-1/2}\kappa_{\infty}^{1/2}(\pi_{\infty}+1).$ (81)
Combining (79), (80), (81) proves the lemma.
B.6 Proof of Theorem 4.5
We decompose $\|h_{\lambda}-\widetilde{h}_{\lambda}\|_{\mathcal{H}}\leq\|h_{\lambda}-\hat{h}_{\lambda}\|_{\mathcal{H}}+\|\hat{h}_{\lambda}-\widetilde{h}_{\lambda}\|_{\mathcal{H}}$ and combine this with (19) and (32) to obtain (34). Similarly, we decompose (6) and use (69) to obtain
${\mathcal{E}}(\widetilde{h}_{\lambda})$ $\leq{\mathcal{E}}(h_{\lambda})+\|J_{\mathbb{P}}h_{\lambda}-J_{\mathbb{P}}\widetilde{h}_{\lambda}\|_{L^{2}_{{\mathbb{P}}}}\leq{\mathcal{E}}(h_{\lambda})+\kappa_{\infty}^{1/2}\|h_{\lambda}-\widetilde{h}_{\lambda}\|_{\mathcal{H}}.$
Combining this with (34) we obtain (35). The second part of the theorem follows from Lemma 2.3 parts (i) and (iii), and because $C_{FS}(\eta,s)$ is increasing in $s$.
B.7 Proof of Theorem 5.1
By assumption we have $\ker J_{\mathbb{P}}^{\ast}=\{0\}$. The equivalence (38) and (39) now follows from (8).
(i): This follows from Theorem 4.5, (34) and (39), and because (31) implies that the norm $\|\widetilde{h}_{\lambda}\|_{\mathcal{H}}$ is given by the left hand side of (40).
(ii): Hypothesis (38) implies $J_{\mathbb{Q}}^{\ast}1=J_{\mathbb{P}}^{\ast}p_{\star}$, by Lemma 2.2. Hence Lemma B.1 (ii) implies that, asymptotically for large $n$,
$u_{\lambda}=n^{-1/2}\big{(}S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{p}_{\star}\big{)}\sim{\mathcal{N}}(0,Q_{\lambda})$ (82)
is normally distributed, where the covariance operator (18) simplifies to
$Q_{\lambda}=J_{\mathbb{Q}}^{\ast}J_{\mathbb{Q}}-(J_{\mathbb{Q}}^{\ast}1)\otimes(J_{\mathbb{Q}}^{\ast}1)+J_{\mathbb{P}}^{\ast}\operatorname{diag}(p_{\star})^{2}J_{\mathbb{P}}-(J_{\mathbb{P}}^{\ast}p_{\star})\otimes(J_{\mathbb{P}}^{\ast}p_{\star}).$
We estimate $Q_{\lambda}$ by its sample analogue given by
$\hat{Q}_{\lambda}=n^{-1}S_{\mathbb{Q}}^{\ast}S_{\mathbb{Q}}-(n^{-1}S_{\mathbb{Q}}^{\ast}\bm{1})\otimes(n^{-1}S_{\mathbb{Q}}^{\ast}\bm{1})+n^{-1}S_{\mathbb{P}}^{\ast}\operatorname{diag}(\bm{p}_{\star})^{2}S_{\mathbb{P}}-(n^{-1}S_{\mathbb{P}}^{\ast}\bm{p}_{\star})\otimes(n^{-1}S_{\mathbb{P}}^{\ast}\bm{p}_{\star}).$
To reduce the dimension, we project the sample variable $u_{\lambda}\mapsto P_{\Pi}u_{\lambda}\eqqcolon\bm{\psi}^{\top}v_{\lambda}$ on the subspace ${\mathcal{H}}_{\Pi}$, with coordinate vector $v_{\lambda}=\langle\bm{\psi},u_{\lambda}\rangle_{\mathcal{H}}$, see (28). Using (72)–(74), we obtain that $v_{\lambda}$ is given by (41) and normally distributed with mean zero and $m\times m$-covariance matrix (42).
(iii): This follows directly from (ii).
B.8 Proof of Lemma 5.3
Denote by $f_{X}(x)=\int_{\mathcal{Y}}f(x,y)\mu_{\mathcal{Y}}(dy)$ and $f_{Y}(x)=\int_{\mathcal{X}}f(x,y)\mu_{\mathcal{X}}(dx)$ the marginal densities, such that ${\mathbb{P}}_{X}(dx)=f_{X}(x)\mu_{\mathcal{X}}(dx)$ and ${\mathbb{P}}_{Y}(dy)=f_{Y}(y)\mu_{\mathcal{Y}}(dy)$. Define their positivity sets $S_{X}\coloneqq\{f_{X}>0\}$ and $S_{Y}\coloneqq\{f_{Y}>0\}$. We claim that ${\mathbb{P}}_{(X,Y)}[S_{X}\times S_{Y}]=1$. Indeed, let $B$ be any measurable set contained in the complement ${\mathcal{Z}}\setminus(S_{X}\times S_{Y})$. Then ${\mathbb{P}}_{(X,Y)}[B]\leq{\mathbb{P}}_{(X,Y)}[B\cap(S_{X}^{c}\times{\mathcal{Y}})]+{\mathbb{P}}_{(X,Y)}[B\cap({\mathcal{X}}\times S_{Y}^{c})]\leq{\mathbb{P}}_{(X,Y)}[S_{X}^{c}\times{\mathcal{Y}}]+{\mathbb{P}}_{(X,Y)}[{\mathcal{X}}\times S_{Y}^{c}]={\mathbb{P}}_{X}[S_{X}^{c}]+{\mathbb{P}}_{Y}[S_{Y}^{c}]=0$, which proves the claim. Hence we can replace $f$ in (53) by $f1_{S_{X}\times S_{Y}}$, where $1_{B}$ denotes the indicator function of a set $B$. By definition we have $01_{S_{X}\times S_{Y}}(x,y)=$ 0 for any $(x,y)\in{\mathcal{Z}}$ such that $f_{X}(x)f_{Y}(y)=0$. As on the other hand we have $({\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y})(dx,dy)=f_{X}(x)f_{Y}(y)(\mu_{X}\otimes\mu_{Y})(dx,dy)$, this proves (44).
Appendix C Independence Test Distributions
This appendix contains the distributions for the independence tests in Section 6.
  • Independent clouds
    $X=X_{0}+\varepsilon_{X},Y=Y_{0}+\varepsilon_{Y},$
    where $X_{0}$ and $Y_{0}$ take values $\{-1,1\}$ with probability 1/2, and $\varepsilon_{X}$ and $\varepsilon_{Y}$ are i.i.d. standard normal.
  • W
    $X\sim\text{Unif}(-1,1),Y=C(X^{2}-0.5)^{2}+\varepsilon,$
    where where $\varepsilon\sim{\mathcal{U}}(0,1)$.
  • Diamond
    $U_{1}=U\cos\frac{\pi}{4}+V\sin\frac{\pi}{4},V_{1}=-U\cos\frac{\pi}{4}+V\sin\frac{\pi}{4},$
    where $U$, $V\sim{\mathcal{U}}(-1,1)$ are independent uniform random variables. Let $(X,Y)=(U_{1},V_{1})$ if $\varepsilon<C$, and $(X,Y)=(U_{2},V_{2})$ otherwise, where $U_{2}$, $V_{2}\sim{\mathcal{U}}(-1,1)$ and $\varepsilon\sim{\mathcal{U}}(0,1)$ are all i.i.d. random variables.
  • Parabola
    $X\sim{\mathcal{U}}(-1,1),Y=CX^{2}+\varepsilon,$
    where $\varepsilon\sim{\mathcal{U}}(0,1)$.
  • Two parabola
    $X\sim{\mathcal{U}}(-1,1),\,Y=(CX^{2}+\varepsilon)V,$
    where $V$ takes values $\{-1,1\}$ with probabilities $1/2$, and $\varepsilon\sim{\mathcal{U}}(0,1)$.
  • Circle
    $X=C\sin(2\pi U)+\varepsilon_{1},\,Y=4.2\cos(2\pi U)+\varepsilon_{2},$
    where $U\sim{\mathcal{U}}(-1,1)$, $\varepsilon_{1}$ and $\varepsilon_{2}$ are i.i.d. standard normal.
  • Variance
    $Y=\varepsilon\sqrt{CX^{2}+1},$
    where $X$ and $\varepsilon$ are i.i.d. standard standard normal.
  • Log
    $Y=C\log X^{2}+\varepsilon,$
    where $X$ and $\varepsilon$ are i.i.d. standard standard normal.

### v2

@@
 Abstract
-We introduce kernel density machines (KDM), a novel density ratio estimator in a reproducing kernel Hilbert space setting.
-KDM applies to general probability measures on countably generated measurable spaces without restrictive assumptions on continuity, or the existence of a Lebesgue density.
+We introduce kernel density machines (KDM), a nonparametric estimator of a Radon–Nikodym derivative, based on reproducing kernel Hilbert spaces.
+KDM applies to general probability measures on countably generated measurable spaces under minimal assumptions.
 For computational efficiency, we incorporate a low-rank approximation with precisely controlled error that grants scalability to large-sample settings.
@@
 Empirical results based on simulated and real data demonstrate the efficacy and precision of KDM.
-Keywords: density ratio estimation, reproducing kernel Hilbert space (RKHS), low-rank approximation, finite-sample guarantees
+Keywords: Radon–Nikodym derivative estimation, reproducing kernel Hilbert space (RKHS), low-rank approximation, finite-sample guarantees
 1 Introduction
-Estimating the density ratio $g_{\star}\coloneqq\frac{d\mathbb{Q}}{d\mathbb{P}}$ of the probability measure $\mathbb{Q}$ with respect to $\mathbb{P}$ is central to many tasks in machine learning, statistics, and applied sciences, such as domain adaption and transfer learning (Pan and Yang, 2010), anomaly detection (Cover and Hart, 1967), causal inference (Imbens and Rubin, 2015), conditional distribution estimation (Racine, 2008; Muandet et al., 2017), and generative modeling.
-Extant approaches to density ratio estimation often assume that both $\mathbb{P}$ and $\mathbb{Q}$ are continuous with respect to the Lebesgue measure, restricting their applicability to fully continuous data.
+Estimating the Radon–Nikodym derivative $g_{\star}\coloneqq\frac{d\mathbb{Q}}{d\mathbb{P}}$ of the probability measure $\mathbb{Q}$ with respect to $\mathbb{P}$ is central to many tasks in machine learning, statistics, and applied sciences, such as domain adaption and transfer learning (Pan and Yang, 2010), anomaly detection (Cover and Hart, 1967), causal inference (Imbens and Rubin, 2015), conditional distribution estimation (Racine, 2008; Muandet et al., 2017), and generative modeling.
+We refer to $g_{\star}$ as the density of ${\mathbb{Q}}$ relative to ${\mathbb{P}}$ throughout, following standard usage in probability theory.
+[^4] This terminology follows the convention of (Jacod and Shiryaev, 1987, Chapter III.3).
+Extant approaches often assume that both $\mathbb{P}$ and $\mathbb{Q}$ admit Lebesgue densities, in which case $g_{\star}$ is commonly referred to as the density ratio, thereby limiting their applicability to purely continuous distributions.
 In contrast, we introduce a kernel-based method that applies to general probability measures on countably generated measurable spaces, encompassing settings with discrete, continuous, mixed, or even singular distributions: kernel density machines (KDM).
-The parsimony of KDM with respect to the underlying assumptions significantly expands the applicability of density ratio estimation.
+The parsimony of KDM with respect to the underlying assumptions significantly expands its versatility with respect to applications.
 We model $g_{\star}$ as the sum of an exogenous prior plus a function in a reproducing kernel Hilbert space (RKHS), which, while possessing attractive theoretical properties, is known to pose computational problems with large data sets.
 To ensure scalability in the sample size and also higher-dimensional applications, we therefore introduce a low-rank approximation that retains critical information while reducing computational costs.
-The approximation error associated with the low-rank density ratio is sharply controlled by the error of a low-rank approximation of the kernel matrix pertaining to the RKHS and the data points.
+The approximation error associated with the low-rank estimator is sharply controlled by the error of a low-rank approximation of the kernel matrix pertaining to the RKHS and the data points.
 This combination of flexibility and computational efficiency makes our method applicable to various real-world scenarios, from structured data in financial modeling, to complex conditional distributions in scientific research.
@@
 These guarantees make KDM reliable across different sample sizes and data configurations.
-As an application, we obtain statistical tests of the prior density ratio and recover the kernel two-sample test of Gretton et al.
-(2012) as a special case.
+As an application, we recover the kernel two-sample test of Gretton et al.
+(2012) as a special case of our framework.
 As a second application, we obtain a novel, consistent estimator of conditional distributions.
@@
 The empirical results suggest KDM to be useful, in particular for larger data sets and higher dimensions than usually considered in RKHS and local nonparametric kernel density estimators (Li and Racine, 2006).
-Our approach builds on recent advances in kernel-based density ratio estimation, but extends beyond existing methods by applying to general probability measures.
+Our approach builds on recent advances in kernel-based density estimation, but extends beyond existing methods by applying to general probability measures.
 Sugiyama et al.
-(2012, 2008) and Filipović et al.
-(2024) are closest to our approach, but are restricted to Lebesgue densities and do not provide theoretical guarantees, respectively.
-Another related work is Zellinger et al.
-(2023), who provide optimal convergence rates of density ratio estimators under rather strong regularity assumptions (source conditions), as used in inverse problems and learning theory.
+(2012, 2008) and Filipovic et al.
+(2025) are related to our approach, but are either restricted to ratios of Lebesgue densities or do not provide the same theoretical guarantees and low-rank approximation error estimates as we do, respectively.
+Another related work is Nguyen et al.
+(2024); Zellinger et al.
+(2023), who provide optimal convergence rates under rather strong regularity assumptions (source conditions), as used in inverse problems and learning theory.
 This is in contrast to our more agnostic framework.
 They also do not discuss computational scalability, as we do with low-rank approximations.
-Another strand of literature, including (Song et al., 2008; Klebanov et al., 2021; Park and Muandet, 2020), estimates conditional expectations directly via conditional mean embeddings (CME).
-This is a powerful and empirically widely used technique.
-But there is no obvious way how to recover density ratio estimates from CMEs.
+Another strand of literature, including (Song et al., 2008; Klebanov et al., 2021; Park and Muandet, 2020), estimates conditional expectations directly via the conditional mean embedding (CME).
+This is a powerful and empirically widely used technique, but there is no obvious way how to recover the density from the CME.
 The remainder of this paper is organized as follows.
-Section 2 formalizes the density ratio estimation problem for general probability measures and introduces our RKHS-based estimator.
+Section 2 formalizes the problem of learning the density for general probability measures and introduces our RKHS-based estimator.
 Section 3 presents theoretical guarantees, including consistency and finite-sample bounds.
@@
 The appendix contains additional results and all proofs.
-2 Density Ratio Learning Problem
+2 General Density Estimation Problem
 Let ${\mathbb{P}}$ and ${\mathbb{Q}}$ be probability measures on some countably generated [^6] Assuming a countably generated $\sigma$-algebra implies that the $L^{2}$ spaces studied below are separable.
@@
 ${\mathbb{Q}}\ll{\mathbb{P}}.$ (1)
-We denote the density ratio (Radon–Nikodym derivative) by $g_{\star}\coloneqq\frac{d{\mathbb{Q}}}{d{\mathbb{P}}}$.
-The goal of this paper is to learn the true density ratio $g_{\star}$ under minimal assumptions from samples of ${\mathbb{P}}$ and ${\mathbb{Q}}$. To this end, we assume that
+We denote the density by $g_{\star}\coloneqq\frac{d{\mathbb{Q}}}{d{\mathbb{P}}}$.
+The goal of this paper is to learn the true density $g_{\star}$ under minimal assumptions from samples of ${\mathbb{P}}$ and ${\mathbb{Q}}$. To this end, we assume that
 $g_{\star}\in L^{2}_{{\mathbb{P}}},$ (2)
@@
 $J_{\mathbb{M}}^{\ast}f=\int_{{\mathcal{Z}}}k(\cdot,z)f(z)\,{\mathbb{M}}(dz),\quad f\in L^{2}_{{\mathbb{M}}},$
-see, e.g., Steinwart and Scovel (2012, Section 2). Given (2), we can represent the hypothesis density ratio as
+see, e.g., Steinwart and Scovel (2012, Section 2). Given (2), we can represent the hypothesis density as
 $g=p_{\star}+J_{\mathbb{P}}h,\quad h\in{\mathcal{H}},$ (4)
@@
 Hence, $p_{\star}=0$ is a possible choice.
-Notably, our specification of the hypothesis density ratio (4) is novel to the literature, as it is more flexible and requires only the milder assumption (2), in contrast to the stronger assumptions typically employed in previous work.
+Notably, our specification of the hypothesis density (4) is novel to the literature, as it is more flexible and requires only the milder assumption (2), in contrast to the stronger assumptions typically employed in previous work.
 For a more detailed comparison, see Section 5.2.3 below.
@@
 In general, the right hand side of (8) is thus not well defined for $\lambda=0$.
-The error function ${\mathcal{E}}(h)$, and the expression in (8), contain the unobservable population density ratio $g_{\star}$ and therefore are not readily available for estimations based on samples of ${\mathbb{P}}$ and ${\mathbb{Q}}$.
+The error function ${\mathcal{E}}(h)$, and the expression in (8), contain the unobservable population density $g_{\star}$ and therefore are not readily available for estimations based on samples of ${\mathbb{P}}$ and ${\mathbb{Q}}$.
 We solve this problem by using the identity $J_{\mathbb{P}}^{\ast}g_{\star}=J_{\mathbb{Q}}^{\ast}1$ justified by the following lemma.
@@
 $h_{\lambda}=(J_{\mathbb{P}}^{\ast}J_{\mathbb{P}}+\lambda)^{-1}(J_{\mathbb{Q}}^{\ast}1-J_{\mathbb{P}}^{\ast}p_{\star}).$ (10)
-To assess how close $p_{\star}+J_{\mathbb{P}}h_{\lambda}$ is to the true density ratio $g_{\star}$, we consider the orthogonal decomposition of the squared population error
+To assess how close $p_{\star}+J_{\mathbb{P}}h_{\lambda}$ is to the true density $g_{\star}$, we consider the orthogonal decomposition of the squared population error
 ${\mathcal{E}}(h_{\lambda})^{2}=\|g_{\star}-p_{\star}-J_{\mathbb{P}}h_{\lambda}\|_{L^{2}_{{\mathbb{P}}}}^{2}=\underbrace{\|g_{\star}-g_{0}\|_{L^{2}_{{\mathbb{P}}}}^{2}}_{\text{projection error}}+\underbrace{\|g_{0}-p_{\star}-J_{\mathbb{P}}h_{\lambda}\|_{L^{2}_{{\mathbb{P}}}}^{2}}_{\text{regularization error}},$ (11)
@@
 Output: rank $m\coloneqq i$, pivot set $\Pi\coloneqq\Pi^{(m)}$, matrices $\bm{L}\coloneqq\bm{L}^{(m)}$, $\bm{R}\coloneqq\bm{B}^{(m)}_{\Pi,:}$
-The following classic result asserts that Algorithm 1 works, see, e.g., Horn and Zhang (2005) or Harbrecht et al.
+The following classic result asserts that Algorithm 1 works, see, e.g., Horn and Zhang (2005), Harbrecht et al.
+(2012) or Filipovic et al.
 (2025).
@@
 5.1 Hypothesis Testing
-As a first application, we obtain statistical tests of the hypothesis that the true equals the prior density ratio,
+As a first application, we obtain statistical tests of the hypothesis that the true equals the prior density,
 […]
@@
 (2017).
-Then the sample variable $u_{\lambda}$ in (82) is the scaled difference $S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{1}$ between the sample kernel mean embeddings of ${\mathbb{Q}}$ and ${\mathbb{P}}$.
+Then the sample variable $u_{\lambda}$ in (82) in the proof of Theorem 5.1 is the scaled difference $S_{\mathbb{Q}}^{\ast}\bm{1}-S_{\mathbb{P}}^{\ast}\bm{1}$ between the sample kernel mean embeddings of ${\mathbb{Q}}$ and ${\mathbb{P}}$.
 As a corollary of Theorem 5.1 we thus recover the kernel two-sample test of Gretton et al.
@@
 and ${\mathbb{P}}_{X\mid Y=y}(dx)=g_{\star}(x,y){\mathbb{P}}_{X}(dx)$, respectively. In the following, we will concentrate on the former. By symmetry, all statements also apply to the latter.
-For any hypothesis density ratio (4) with $h\in{\mathcal{H}}$, we thus obtain the corresponding model of the true conditional expectation ${\mathbb{E}}[f(Y)\mid X=x]=\int_{\mathcal{Y}}f(y){\mathbb{P}}_{Y\mid X=x}(dy)$ given by
+For any hypothesis density (4) with $h\in{\mathcal{H}}$, we thus obtain the corresponding model of the true conditional expectation ${\mathbb{E}}[f(Y)\mid X=x]=\int_{\mathcal{Y}}f(y){\mathbb{P}}_{Y\mid X=x}(dy)$ given by
 ${\mathbb{E}}_{h}[f(Y)\mid X=x]\coloneqq\int_{\mathcal{Y}}f(y)(p_{\star}(x,y)+h(x,y))\,{\mathbb{P}}_{Y}(dy),\quad\text{for $f\in L^{2}_{{\mathbb{P}}_{Y}}$.}$ (46)
@@
 The last approximation in (51) is practically motivated such that we integrate $f$ with respect to a bona-fide conditional distribution.
-[^12] The density ratio $g_{\star}$ satisfies the implicit structural properties $g_{\star}(x,y)\geq 0$ ${\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}$-a.s., $\int_{\mathcal{Y}}g_{\star}(x,y){\mathbb{P}}_{Y}(dy)=1$ ${\mathbb{P}}_{X}$-a.s., and $\int_{\mathcal{X}}g_{\star}(x,y){\mathbb{P}}_{X}(dx)=1$ ${\mathbb{P}}_{Y}$-a.s.
+[^13] The density $g_{\star}$ satisfies the implicit structural properties $g_{\star}(x,y)\geq 0$ ${\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}$-a.s., $\int_{\mathcal{Y}}g_{\star}(x,y){\mathbb{P}}_{Y}(dy)=1$ ${\mathbb{P}}_{X}$-a.s., and $\int_{\mathcal{X}}g_{\star}(x,y){\mathbb{P}}_{X}(dx)=1$ ${\mathbb{P}}_{Y}$-a.s.
 The estimator $\widetilde{h}_{\lambda}$ in (49) requires i.i.d.
-samples $z_{{\mathbb{P}},1},\dots,z_{{\mathbb{P}},n}$ of ${\mathbb{P}}\coloneqq{\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}$ and $z_{{\mathbb{Q}},1},\dots,z_{{\mathbb{Q}},n}$ of ${\mathbb{Q}}\coloneqq{\mathbb{P}}_{(X,Y)}$.
+samples $z_{{\mathbb{P}},1},\dots,z_{{\mathbb{P}},n}$ of ${\mathbb{P}}={\mathbb{P}}_{X}\otimes{\mathbb{P}}_{Y}$ and $z_{{\mathbb{Q}},1},\dots,z_{{\mathbb{Q}},n}$ of ${\mathbb{Q}}={\mathbb{P}}_{(X,Y)}$.
 If we could only sample from the joint distribution, we would consider an i.i.d.
@@
 Kernel conditional mean embeddings
-In view of the decomposition (48) it is natural to compare conditional expectations (46) calculated from a hypothesis density ratio (4) to the conditional mean embedding introduced in Song et al.
+In view of the decomposition (48) it is natural to compare conditional expectations (46) calculated from a hypothesis density (4) to the conditional mean embedding introduced in Song et al.
 (2009).
@@
 Third, the conditional mean embedding acts through (54) only on functions $f\in{\mathcal{H}}_{\mathcal{Y}}$.
-Our approach, via the density ratio $g_{\star}$, overcomes the above limitations of the traditional conditional mean embedding approach.
+Our approach, via the density $g_{\star}$, overcomes the above limitations of the traditional conditional mean embedding approach.
 First, our construction of $\mu_{Y\mid X=x}$ is more elementary and feasible under verifiable technical assumptions, such as (1) and (2).
@@
 In fact, they assume the assumptions of Lemma 5.3 hold for the Lebesgue measures $\mu_{\mathcal{X}}(dx)=dx$ and $\mu_{\mathcal{Y}}(dy)=dy$.
-In this case, the “density ratio $w(x,y)$” in Suzuki et al.
-(2009) is equal to our density ratio $g_{\star}(x,y)$.
-Their estimator of the density ratio is based on a sampling scheme, similar to Filipović et al.
+In this case, the “density $w(x,y)$” in Suzuki et al.
+(2009) is equal to our density $g_{\star}(x,y)$.
+Their estimator of the density is based on a sampling scheme, similar to Filipovic et al.
 (2025), which is different from ours.
@@
 (2023) for a recent overview.
-In contrast, our approach is nonparametric functional analytic and learns the density ratio in a reproducing kernel Hilbert space, which comes with asymptotic theory and finite sample guarantees.
+In contrast, our approach is nonparametric functional analytic and learns the density in a reproducing kernel Hilbert space, which comes with asymptotic theory and finite sample guarantees.
 Isotonic distribution regression in Henzi et al.
@@
 7 Conclusion
-Kernel density machines (KDM) is a comprehensive data-driven framework for estimating the density ratio (Radon-Nikodym derivative) $\frac{d\mathbb{Q}}{d\mathbb{P}}$ from i.i.d.
+Kernel density machines (KDM) is a comprehensive data-driven framework for estimating the Radon–Nikodym derivative (density) $\frac{d\mathbb{Q}}{d\mathbb{P}}$ from i.i.d.
 samples of $\mathbb{P}$ and $\mathbb{Q}$ in a reproducing kernel Hilbert hypothesis space.